### v1

Abstract
The evolution of diffusion models has greatly impacted video generation and understanding. Particularly, text-to-video diffusion models (VDMs) have significantly facilitated the customization of input video with target appearance, motion, etc. Despite these advances, challenges persist in accurately distilling motion information from video frames. While existing works leverage the consecutive frame residual as the target motion vector [16], they inherently lack global motion context and are vulnerable to frame-wise distortions. To address this, here we present Spectral Motion Alignment (SMA), a novel framework that refines and aligns motion vectors using Fourier and wavelet transforms. SMA learns motion patterns by incorporating frequency-domain regularization, facilitating the learning of whole-frame global motion dynamics, and mitigating spatial artifacts. Extensive experiments demonstrate SMA’s efficacy in improving motion transfer while maintaining computational efficiency and compatibility across various video customization frameworks.
Keywords:
Diffusion models Video Motion transfer Wavelet transform
Figure 1: One-shot Video Motion Transfer via Spectral Motion Alignment using Cascaded Video Diffusion. SMA facilitates the capture of long-range (top) and complex (bottom) motion patterns within videos. Visit here for comprehensive view of the videos.
1 Introduction
The progression of diffusion models [31, 12] has played a pivotal role in advancing video generation and comprehension. Notably, recent advancements in text-conditioned Video Diffusion Models (VDMs, [13, 11, 30, 4, 9, 45, 3, 22]) have significantly enhanced the generation of highly realistic and superior-quality videos. Given the multifaceted nature of the video, encompassing motion dynamics, foreground object appearance, etc., several studies [49, 39, 16, 42, 41, 27, 36] aim to disentangle, modify, and compose these signals according to user intent. Additionally, various approaches specifically target to customize the motion, pose, shape, or appearance of real-world videos.
In the context of motion customization, multiple studies explore how to distill motion information from pre-trained video diffusion models. Recent research suggests that motion patterns are inherently encoded in the underlying dependencies between latent frames or epsilon noises. As noted by [49], videos with similar motion tend to exhibit similar connectivity between latent frames. Additionally, [16] utilizes residual vectors between consecutive latent frames as "motion vectors," in line with optical flow principles, where residuals in consecutive frames represent motion dynamics. Specifically, they finetune temporal attention layers of pretrained VDM to align the ground-truth image-space residuals with their predicted denoised estimates. Their work reveals that aligning these predicted and ground-truth motion vectors corresponds to aligning epsilon (noise) residuals, readily obtainable from any off-the-shelf video diffusion models.
Figure 2: Overview. The proposed Spectral Motion Alignment (SMA) framework distills the motion information in frequency-domain. Considering the (latent) frame residuals as motion vectors, we first derive the denoised motion vector estimates. Then, the motion vector $\delta{\boldsymbol{v}}_{0}^{n}$ and its estimate $\delta\hat{{\boldsymbol{v}}}_{0}^{n}$ are aligned in both pixel-domain and frequency-domain. Our regularization includes (1) global motion alignment based on 1D wavelet-transform, and (2) local motion refinement based on 2D Fourier transform.
While these efforts have made technical advances in motion distillation, the pitfalls of using latent frame residuals as motion vectors have not been fully studied. This paper focuses on the potential limitations of current motion estimation in diffusion models: (a) lack of global context information, and (b) vulnerability to frame-wise distortions. Since frame residuals may capture local motion patterns but are blind to whole-frame motion dynamics, for better motion dynamics modeling, we have to understand the whole-frame global context information during motion distillation. Furthermore, while the frame residuals contain motion information, they may also contain inevitable disruptive variations that are unrelated to motion. These variations may include abrupt changes in the background, distortions, consecutive frame inconsistencies, or lighting condition changes.
To address these challenges, here we present Spectral Motion Alignment (SMA), a novel Fourier and wavelet-transform-based motion vector refinement and alignment framework. This includes two primary components: First, to learn the global motion context, we propose the spectral alignment loss between predicted and ground-truth motion vectors within the wavelet domain. This facilitates the learning of multi-scale motion dynamics by leveraging rich wavelet-domain representations of video considering the global frame transitions. Moreover, to mitigate the spatial artifacts and inconsistency in motion vectors, we propose 2D FFT-based motion vector refinement that aligns the amplitude and phase spectrum of ground truth and predicted motion vectors with prioritizing low-frequency components. This is because the high-frequency components in motion vectors may be associated with frame-wise non-motion-related artifacts (Figure 6). To sum up, we encourage accurate motion transfer via harmonized global and local levels of frequency-domain regularization.
Our contributions are summarized as follows:
  • We introduce the Spectral Motion Alignment (SMA), a frequency-domain motion alignment framework that learns the underlying motion dynamics of input video via frequency-based regularization. Since conventional methods mostly rely on pixel-domain motion distillation, the proposed framework is orthogonal to existing motion customization models.
  • SMA imposes negligible memory and computational burdens, as most off-the-shelf VDMs can readily compute estimates of motion vectors. For instance, VMC [16] with SMA demonstrates lightweight (15GB vRAM) and rapid (< 5 minutes) training.
  • We validate the advantage of SMA across diverse motion patterns and subjects, and across various video motion transfer frameworks such as Video Diffusion-based [49], Cascaded Video Diffusion-based [16], Text-to-Image Diffusion-based [38], and ControlNet-based models [7].
2 Preliminaries
2.1 Diffusion Models.
Diffusion models [31, 12] generate samples from the Gaussian noise through reverse denoising processes. Given a clean sample ${\boldsymbol{x}}_{0}\sim p_{\operatorname{data}}({\boldsymbol{x}})$, the forward process is defined as a Markov chain with forward conditional densities
$\begin{split}p({\boldsymbol{x}}_{t}\>|\>{\boldsymbol{x}}_{t-1})={\mathcal{N}}({\boldsymbol{x}}_{t}\>|\>\beta_{t}{\boldsymbol{x}}_{t-1},(1-\beta_{t})I)\\ p_{t}({\boldsymbol{x}}_{t}\>|\>{\boldsymbol{x}}_{0})={\mathcal{N}}({\boldsymbol{x}}_{t}\>|\>\sqrt{{\bar{\alpha}}}{\boldsymbol{x}}_{0},(1-{\bar{\alpha}})I),\end{split}$ (1)
where ${\boldsymbol{x}}_{t}\in\mathbb{R}^{d}$ is a noisy latent variable at time $t$ that has the same dimension as ${\boldsymbol{x}}_{0}$, and $\beta_{t}$ denotes an increasing sequence of noise schedule where $\alpha_{t}\coloneqq 1-\beta_{t}$ and ${\bar{\alpha}}_{t}\coloneqq\Pi_{i=1}^{t}\alpha_{i}$. Then, the goal of diffusion model training is to obtain a residual denoiser ${\boldsymbol{\epsilon}}_{\theta^{*}}$:
$\theta^{*}:=\mathop{\mathrm{argmin}}_{\theta}{\mathbb{E}}_{{\boldsymbol{x}}_{t}\sim p_{t}({\boldsymbol{x}}_{t}\>|\>{\boldsymbol{x}}_{0}),{\boldsymbol{x}}_{0}\sim p_{\operatorname{data}}({\boldsymbol{x}}_{0}),{\boldsymbol{\epsilon}}\sim{\mathcal{N}}(0,I)}\big{[}\left\lVert{\boldsymbol{\epsilon}}_{\theta}({\boldsymbol{x}}_{t},t)-{\boldsymbol{\epsilon}}\right\rVert\big{]}.$ (2)
The reverse sampling from $q({\boldsymbol{x}}_{t-1}|{\boldsymbol{x}}_{t},{\boldsymbol{\epsilon}}_{\theta^{*}}({\boldsymbol{x}}_{t},t))$ is then achieved by
${\boldsymbol{x}}_{t-1}=\frac{1}{\sqrt{\alpha_{t}}}\Big{(}{\boldsymbol{x}}_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\boldsymbol{\epsilon}_{\theta^{*}}({\boldsymbol{x}}_{t},t)\Big{)}+\tilde{\beta}_{t}{\boldsymbol{\epsilon}},$ (3)
where ${\boldsymbol{\epsilon}}\sim{\mathcal{N}}(0,{\boldsymbol{I}})$ and $\tilde{\beta}_{t}\coloneqq\frac{1-{\bar{\alpha}}_{t-1}}{1-{\bar{\alpha}}_{t}}\beta_{t}$. To accelerate sampling, DDIM [32] further proposes another sampling method as follows:
${\boldsymbol{x}}_{t-1}=\sqrt{{\bar{\alpha}}_{t-1}}\hat{{\boldsymbol{x}}}_{0}(t)+\sqrt{1-{\bar{\alpha}}_{t-1}-\eta^{2}\tilde{\beta_{t}}^{2}}{\boldsymbol{\epsilon}}_{\theta^{*}}({\boldsymbol{x}}_{t},t)+\eta\tilde{\beta}_{t}{\boldsymbol{\epsilon}},$ (4)
where $\eta\in[0,1]$ is a stochasticity parameter, and $\hat{{\boldsymbol{x}}}_{0}(t)$ is the denoised estimate which can be equivalently derived using Tweedie’s formula [8]:
$\hat{{\boldsymbol{x}}}_{0}(t)\coloneqq\frac{1}{\sqrt{{\bar{\alpha}}_{t}}}({\boldsymbol{x}}_{t}-\sqrt{1-{\bar{\alpha}}_{t}}{\boldsymbol{\epsilon}}_{\theta^{*}}({\boldsymbol{x}}_{t},t)).$ (5)
For a text-guided generation, diffusion models are often trained with the textual embedding $c$. Throughout this paper, we will often omit $c$ from ${\boldsymbol{\epsilon}}_{\theta}({\boldsymbol{x}}_{t},t,c)$ if it does not lead to notational ambiguity.
2.1.1 Video Diffusion Models.
Video diffusion models [13, 11, 45] further attempt to model the video data distribution. Specifically, Let $({\boldsymbol{v}}^{n})_{n\in\{1,\dots,N\}}$ represents the $N$-frame input video sequence. Then, for a given $n$-th frame ${\boldsymbol{v}}^{n}\in\mathbb{R}^{d}$, let ${\boldsymbol{v}}^{1:N}\in\mathbb{R}^{N\times d}$ represents a whole video vector. Let $\boldsymbol{v}_{t}^{n}=\sqrt{\bar{\alpha}_{t}}\boldsymbol{v}^{n}+\sqrt{1-\bar{\alpha}_{t}}\boldsymbol{\epsilon}_{t}^{n}$ represents the $n$-th noisy frame latent sampled from $p_{t}(\boldsymbol{v}_{t}^{n}|\boldsymbol{v}^{n})$, where $\boldsymbol{\epsilon}_{t}^{n}\sim\mathcal{N}(0,I)$. We similarly define $({\boldsymbol{v}}_{t}^{n})_{n\in 1,\dots,N}$, ${\boldsymbol{v}}_{t}^{1:N}$, and ${\boldsymbol{\epsilon}}^{1:N}$. The goal of video diffusion model training is then to obtain a residual denoiser ${\boldsymbol{\epsilon}}_{\theta}$ with textual condition $c$ and video input that satisfies:
$\min_{\theta}{\mathbb{E}}_{{\boldsymbol{v}}_{t}^{1:N},{\boldsymbol{v}}^{1:N},{\boldsymbol{\epsilon}}^{1:N},c}\big{[}\left\lVert{\boldsymbol{\epsilon}}_{\theta}({\boldsymbol{v}}_{t}^{1:N},t,c)-{\boldsymbol{\epsilon}}^{1:N}\right\rVert\big{]},$ (6)
where ${\boldsymbol{\epsilon}}_{\theta}({\boldsymbol{v}}_{t}^{1:N},t,c),{\boldsymbol{\epsilon}}^{1:N}\in\mathbb{R}^{N\times d}$. In this work, we denote the predicted noise of $n$-th frame as ${\boldsymbol{\epsilon}}_{\theta}^{n}({\boldsymbol{v}}_{t}^{1:N},t,c)\in\mathbb{R}^{d}$.
2.2 Fourier and Wavelet Analysis
Spectral analysis techniques transform time-domain or pixel-domain signals (such as video frames) into the frequency domain, revealing the frequency components and their intensities. This transformation is valuable in video understanding as it helps in identifying repetitive motion patterns and underlying structures that may not be visible in the time or pixel domain.
Fourier Transform. Let ${\boldsymbol{v}}^{n}\in\mathbb{R}^{H\times W}$ represents the $n$-th 2D video frame. Then, its frequency spectrum at coordinate $(a,b)$ is given as follows:
$\mathcal{F}_{{\boldsymbol{v}}^{n}}(a,b)=\sum_{x=0}^{H-1}\sum_{y=0}^{W-1}{\boldsymbol{v}}^{n}(x,y)e^{-i2\pi(\frac{ax}{H}+\frac{by}{W})},$ (7)
where ${\boldsymbol{v}}^{n}(x,y)$ means the pixel value at coordinate $(x,y)$. The output frequency spectrum is represented as $\mathcal{F}_{{\boldsymbol{v}}^{n}}(a,b)=R(a,b)+I(a,b)i$, where $R(a,b),I(a,b)\in\mathbb{R}$ represents real and imaginary part, respectively. Then, the amplitude and phase is derived as follows:
$|\mathcal{F}_{{\boldsymbol{v}}^{n}}(a,b)|=\sqrt{R(a,b)^{2}+I(a,b)^{2}},\angle\mathcal{F}_{{\boldsymbol{v}}^{n}}(a,b)=\arctan\Big{(}\frac{I(a,b)}{R(a,b)}\Big{)}.$ (8)
Here, amplitude often indicates the existence of corresponding spatial frequencies within the image, often characterized by edges, textures, patterns, etc. Phase often encodes the structural spatial relationship.
Wavelet Transform. We first provide a brief introduction to the frame theory. Frames in signal processing are an extension of basis sets but with a relaxation that allows for redundancy. Let $\Psi=[\psi_{1}\dots\psi_{m}]\in\mathbb{R}^{n\times m}$, where $\{\psi_{k}\}_{k=1}^{m}$ is a set of functions in a Hilbert space $H$. Then, $\{\psi_{k}\}_{k=1}^{m}$ is called as a frame if it satisfies the following inequality:
$\alpha\|f\|^{2}\leq\|\langle f,\Psi\rangle\|^{2}\leq\beta\|f\|^{2},\forall f\in H$ (9)
where $c_{k}\coloneqq\langle f,\psi_{k}\rangle$ is a $k$-th frame expansion coefficient, and $\alpha,\beta$ are frame bounds. The original signal $f$ can be exactly reconstructed from expansion coefficients using the dual frame operator $\tilde{\Psi}$ which satisfies the so-called frame condition: $\tilde{\Psi}\Psi^{T}=I$.
Wavelet frames, renowned for capturing multi-resolution scale features, are among the most prevalently utilized frame representations in signal processing. Let $\psi(t)$ represent a mother wavelet that can be shifted and scaled. For a function ${\boldsymbol{v}}(t)\in L^{2}(\mathbb{R})$, the wavelet transform can be expressed as:
[…]
which serves as an expansion coefficient. In the case of discrete wavelet transform (DWT), it uses a finite set of wavelet and scaling functions derived from a chosen wavelet family. Specifically, the mother wavelet is shifted and scaled by powers of two as follows:
$\psi_{j,k}(t)=\frac{1}{\sqrt{2^{j}}}\psi(2^{-j}t-k).$ (11)
Then, the DWT of a signal ${\boldsymbol{v}}[n]$ is given by:
$\mathcal{W}_{{\boldsymbol{v}}}(j,k)=\langle{\boldsymbol{v}}(t),\psi_{j,k}(t)\rangle.$ (12)
Then, the original signal can be recovered from inverse DWT. In practice, this discrete wavelet transform can be implemented by convolution using an appropriate choice of filter bank.
3 Spectral Motion Alignment
Given an input video, our main goal is to develop a novel frequency-domain motion alignment framework that (a) distills the motion patterns $M^{*}$ of input video, and (b) transfers the motion patterns $M^{*}$ to output video within varied contexts, e.g. Cars with motion $M^{*}$rightarrow Tanks with motion $M^{*}$.
Conventional motion distillation methods mostly rely on pixel-domain motion distillation objectives. However, frequency-domain analysis can further capture underlying motion patterns across a spectrum of frequency levels that mainly constitute motion. Accordingly, we propose Spectral Motion Alignment (SMA), a novel Fourier and wavelet-transform-based motion vector refinement and alignment framework. More details follow.
3.1 Denoised Motion Vector Estimation
To distill the motion information, we first estimate the motion vector of the input video following [16]. The intuition is that residual vectors between consecutive frames may include information about the motion trajectories. Define the $n$-th frame residual vector, namely motion vector at time $t\geq 0$ as
$\delta{\boldsymbol{v}}_{t}^{n}\coloneqq{\boldsymbol{v}}_{t}^{n+1}-{\boldsymbol{v}}_{t}^{n},$ (13)
where the epsilon residual vector $\delta{\boldsymbol{\epsilon}}_{t}^{n}$ is similarly defined. This $\delta{\boldsymbol{v}}_{t}^{n}$ can be acquired through the following diffusion kernel [16]:
$p(\delta{\boldsymbol{v}}_{t}^{n}\>|\>\delta{\boldsymbol{v}}_{0}^{n})={\mathcal{N}}(\delta{\boldsymbol{v}}_{t}^{n}\>|\>\sqrt{{\bar{\alpha}}_{t}}\delta{\boldsymbol{v}}_{0}^{n},2(1-{\bar{\alpha}}_{t})I).$ (14)
Moreover, [16] shows that the ground-truth motion vector in pixel space $\delta{\boldsymbol{v}}_{0}^{n}$ can be derived as follows:
$\delta{\boldsymbol{v}}_{0}^{n}=\frac{1}{\sqrt{{\bar{\alpha}}_{t}}}\Big{(}\delta{\boldsymbol{v}}_{t}^{n}-\sqrt{1-{\bar{\alpha}}_{t}}\delta{\boldsymbol{\epsilon}}_{t}^{n}\Big{)}.$ (15)
Similarly, one can obtain the denoised motion vector estimate $\delta\hat{{\boldsymbol{v}}}_{0}^{n}$ by using Tweedie’s formula as follows:
$\hat{{\boldsymbol{v}}}_{0}^{1:N}(t)\coloneqq\frac{1}{\sqrt{{\bar{\alpha}}_{t}}}\big{(}{\boldsymbol{v}}_{t}^{1:N}-\sqrt{1-{\bar{\alpha}}_{t}}{\boldsymbol{\epsilon}}_{\theta}({\boldsymbol{v}}_{t}^{1:N},t)\big{)},$ (16)
where $\hat{{\boldsymbol{v}}}_{0}^{1:N}(t)$ is an empirical Bayes optimal posterior expectation $\mathbb{E}[{\boldsymbol{v}}_{0}^{1:N}\>|\>{\boldsymbol{v}}_{t}^{1:N}]$.
In the context of motion transfer, the authors in [16] finetune the temporal attention layers of VDM by aligning each motion vector $\delta{\boldsymbol{v}}_{0}^{n}$ and its denoised estimate $\delta\hat{{\boldsymbol{v}}}_{0}^{n}(t)$:
$\min_{\theta}\mathbb{E}_{t,n,{\boldsymbol{\epsilon}}^{t,n},{\boldsymbol{\epsilon}}^{t,n+1}}\Big{[}\ell_{\text{align}}\big{(}\delta{\boldsymbol{v}}_{0}^{n},\delta\hat{{\boldsymbol{v}}}_{0}^{n}(t)\big{)}\Big{]}.$ (17)
While these advancements in motion distillation mark significant progress, Figure 1, 4 and 6 indicate that [16] still has potential for further refinement.
3.2 Spectral Global Motion Alignment
One of the primary limitations in (17) is that it may not fully encapsulate the global motion dynamics. This is partly because [16] employs ‘frame-wise’ matching losses, such as $\ell_{2}$-distance. While this is effective, it may overlook the comprehensive motion dynamics by focusing solely on pairwise frame comparisons.
To mitigate these problems, we explore the use of wavelet transforms in motion distillation. In this paper, we use Haar wavelet, whose low and high pass filters are given as follows:
[…]
which is implemented using the multi-scale Haar filter bank. Then, given the sequence of motion vectors $\delta{\boldsymbol{v}}_{0}=(\delta{\boldsymbol{v}}_{0}^{n})_{n\in\{1,\dots,N-1\}}$ and its denoised estimates $\delta\hat{{\boldsymbol{v}}}_{0}(t)=\big{(}\delta\hat{{\boldsymbol{v}}}_{0}^{n}(t)\big{)}_{n\in\{1,\dots,N-1\}}$, we consider ($N-1$)-length time-dependent 1D arrays from arbitrary spatial pixel dimension $s\in\{1,\dots d\}$. The corresponding 1D array of motion vector is denoted by $\delta{\boldsymbol{v}}_{0,s}$ and $\delta\hat{{\boldsymbol{v}}}_{0,s}(t)\in\mathbb{R}^{N-1}$ (Figure 2).
Then, the frequency-matching loss between $\delta{\boldsymbol{v}}_{0}$ and $\delta\hat{{\boldsymbol{v}}}_{0}(t)$ is defined with DWT in (12) as follows:
$\ell_{\text{global}}(\delta{\boldsymbol{v}}_{0},\delta\hat{{\boldsymbol{v}}}_{0}(t))=\mathbb{E}_{t,s,j,k}\Big{[}\|{\mathcal{W}}_{\delta{\boldsymbol{v}}_{0,s}}(j,k)-{\mathcal{W}}_{\delta\hat{{\boldsymbol{v}}}_{0,s}(t)}(j,k)\|_{1}\Big{]}.$ (19)
Considering that the wavelet transform allows multi-resolution analysis of motion vectors, it enables us to handle motions at various scales and frequencies effectively. This could be particularly beneficial for complex scenes with varying motion speeds and types, ensuring that subtle motions are captured and transferred more accurately.
3.3 Spectral Local Motion Refinement
Another problem in (17) is that the residuals between consecutive frames, while rich in motion information, may also encapsulate high-frequency local distortions, background noise, and other non-motion-related artifacts. Accordingly, we focus on particularly prioritizing low-to-moderate spatial frequency components. Specifically, following the amplitude and phase spectrum definition in (8), we define amplitude and phase matching loss as follows:
$\begin{split}&\ell_{\text{l}ocal}^{A}(\delta{\boldsymbol{v}}_{0}^{n},\delta\hat{{\boldsymbol{v}}}_{0}^{n}(t))=\mathbb{E}_{t,n,a,b}\Big{[}\omega(a,b)\||\mathcal{F}_{\delta{\boldsymbol{v}}_{0}^{n}}(a,b)|-|\mathcal{F}_{\delta\hat{{\boldsymbol{v}}}_{0}^{n}(t)}(a,b)|\|_{1}\Big{]},\\ &\ell_{\text{l}ocal}^{P}(\delta{\boldsymbol{v}}_{0}^{n},\delta\hat{{\boldsymbol{v}}}_{0}^{n}(t))=\mathbb{E}_{t,n,a,b}\Big{[}\omega(a,b)\|\angle\mathcal{F}_{\delta{\boldsymbol{v}}_{0}^{n}}(a,b)-\angle\mathcal{F}_{\delta\hat{{\boldsymbol{v}}}_{0}^{n}(t)}(a,b)\|_{1}\Big{]},\end{split}$ (20)
where the frequency domain weighting $\omega(a,b)$ is defined as
$\omega(a,b)=\Big{[}\big{(}\frac{H}{2}\big{)}^{2}+\big{(}\frac{W}{2}\big{)}^{2}\Big{]}^{\delta}-\Big{[}\big{(}a-\frac{H}{2}\big{)}^{2}+\big{(}b-\frac{W}{2}\big{)}^{2}\Big{]}^{\delta}+1$ (21)
for $0<a<H,0<b<W$, and otherwise, set to zero. This introduces a weighting [40] that prioritizes low-frequency components for $\delta>0$.
Figure 3: Comparison within MotionDirector framework.
3.4 Inference Pipeline
To sum up, the overall spectral motion alignment framework is given as follows:
$\min_{\theta}\mathbb{E}_{t,n,{\boldsymbol{\epsilon}}_{t}^{n},{\boldsymbol{\epsilon}}_{t}^{n+1}}\Big{[}\ell_{\text{align}}\big{(}\delta{\boldsymbol{v}}_{0}^{n},\delta\hat{{\boldsymbol{v}}}_{0}^{n}(t)\big{)}+\lambda_{g}\ell_{\text{global}}\big{(}\delta{\boldsymbol{v}}_{0},\delta\hat{{\boldsymbol{v}}}_{0}(t)\big{)}+\lambda_{l}\ell_{\text{l}ocal}\big{(}\delta{\boldsymbol{v}}_{0}^{n},\delta\hat{{\boldsymbol{v}}}_{0}^{n}(t)\big{)}\Big{]},$ (22)
[…]
This Spectral Motion Alignment (SMA) is universally adaptable across various motion distillation frameworks. While diverse diffusion-based motion distillation frameworks adopt their pixel-domain motion learning objectives, the proposed frequency-domain alignment seamlessly integrates with these arbitrary objectives. Moreover, different motion distillation frameworks target specific parameters $\theta$ for fine-tuning, varying from temporal attention layers [16] to dual-path LoRAs [49]. We empirically demonstrate global compatibility of the proposed spectral motion alignment with diverse neural architectures and parameterizations. Pseudo-code is provided in the appendix.
Figure 4: Comparison within VMC framework, using Show-1 cascaded model (top) and Zeroscope (bottom) non-cascaded model. Visit our project page for an in-depth view.
4 Experiments using T2V Diffusion Models
4.1 Experimental Setting
To assess the capability of Spectral Motion Alignment (SMA) to capture accurate motion within contemporary diffusion-based motion learning frameworks, we curated a dataset comprising 30 text-video pairs sourced from the publicly available DAVIS [24] and WebVid-10M [2] collections. This dataset is deliberately designed to cover a broad spectrum of motion types and subjects, with video lengths ranging between 8 and 16 frames. For this study, we leverage two foundational text-to-video diffusion models: Zeroscope [33], a non-cascaded VDM, and Show-1 [45], a cascaded VDM. More details are provided in appendix.
4.2 Baselines
4.2.1 MotionDirector.
MotionDirector [49] pioneered the concept of motion customization within video diffusion literature. To distinctively tailor the appearance and motion of a video, [49] developed a unique dual-path framework employing Low-Rank Adaptation (LoRA, [14]). Within this architecture, spatial LoRAs are optimized using the diffusion training loss on a single, randomly chosen frame, rather than on the full video sequence. Conversely, temporal LoRAs are trained by considering all frames of the video, using an anchor frame based appearance-debiased temporal loss.
4.2.2 VMC.
VMC [16] achieves state-of-the-art performance in motion customization through their novel epsilon residual matching objective, facilitating efficient motion distillation within cascaded video diffusion framework. In their work, they showcased the efficient of motion distillation within cascaded video diffusion models, by optimizing only keyframe generation model while freezing later modules. As VMC represents the most closely related work to ours, we conduct a two-fold validation within this baseline, including both cascaded and non-cascaded video diffusion scenarios.
4.2.3 Qualitative Comparison.
Fig. 3 offers a qualitative analysis comparing MotionDirector’s performance with and without the implementation of SMA. Likewise, Fig. 4 examines VMC’s performance with and without SMA implementation. Specifically, videos produced using a cascaded diffusion pipeline are displayed in Fig. 4-top, whereas those generated through a non-cascaded diffusion model are shown in the bottom. Videos generated without SMA correctly capture the appearance as dictated by the target text, but fail to accurately mirror the original video’s motion patterns. In contrast, SMA significantly enhances motion understanding, precisely distinguishing between dynamic and static elements in the scene. For instance, in the last example of Fig. 3, MotionDirector with SMA produces a video where only the eagle moves from right to left, whereas without SMA, the video inaccurately depicts the ground moving alongside the eagle.
4.2.4 Quantitative Comparison.
The results of our quantitative evaluation are presented in Table 1. We employ CLIP encoders [26] for automated evaluation. To evaluate text-video alignment [10], we measure the average cosine similarity between the target text prompt and the frames generated. Regarding frame consistency, we extract CLIP image features for each frame in the output video and subsequently calculate the average cosine similarity among all frame pairs in the video. For human evaluation, we conduct a user study with 42 participants to assess three key aspects, guided by the following questions: (1) Editing Accuracy: Is the output video accurately edited, reflecting the target text? (2) Temporal Consistency: Is the transition between frames smooth and consistent? (3) Motion Accuracy: Is the motion of the input video accurately preserved in the output video? Tab. 1 demonstrates that SMA enhances the performance of MotionDirector and VMC across all measured metrics.
Automatic Metrics User Study
Method Text-Align Temp-Con Edit-Acc Temp-Con Motion-Acc
MotionDirector 0.7550 0.9780 3.19 3.07 2.67
MotionDirector w/ Ours 0.8081 0.9784 4.14 3.89 3.88
VMC (Show-1) 0.8066 0.9742 3.25 3.22 2.72
VMC (Show-1) w/ Ours 0.8193 0.9776 3.85 3.84 3.92
VMC (Zeroscope) 0.8223 0.9560 2.95 2.83 2.47
VMC (Zeroscope) w/ Ours 0.8425 0.9578 4.07 3.84 4.07
Table 1: Quantitative evaluation of SMA within text-to-video based frameworks.
Figure 5: Comparison within Tune-A-Video and ControlVideo (Depth) baseline.
5 Experiments using T2I Video Diffusion Models
5.1 Experimental Setting
To assess the effectiveness of SMA within methodologies built on text-to-image diffusion model, we use the same text-video pairs previously assembled in Sec. 4.1. The resolution for all produced videos is standardized to 512x512. In this experiment, Stable Diffusion v1-5 [28] and ControlNet-Depth [46] are utilized.
5.2 Baselines
5.2.1 Tune-A-Video.
Tune-A-Video [38] was at the forefront of image-diffusion-based video editing at a time when video diffusion models were not yet publicly accessible. It begins with transforming a pretrained T2I diffusion model to psuedo T2V model by adding temporal attention layers and expanding spatial self-attention into spatio-temporal attention. Attention projection layers within this pseudo T2V model are then fine-tuned on a specific input video, utilizing the diffusion training objective (Eq. 2).
5.2.2 ControlVideo.
ControlVideo [48] is another one-shot-based video editing method, starting from pretrained T2I model. ControlVideo extends ControlNet [46] from image to video to incorporate structural cues obtained from the input video. Subsequently, akin to Tune-A-Video [38], attention projection layers of the backbone U-Net and ControlNet are customized on the input video.
5.3 Qualitative Comparison.
Fig. 5-(top) demonstrates the qualitative benefits of incorporating SMA into the Tune-A-Video method. Without SMA, Tune-A-Video often experiences flickering in foreground objects, a problem that SMA significantly reduces. Fig. 5-(bottom) illustrates the qualitative improvements brought by integrating SMA with the ControlVideo framework. In ControlVideo, depth control aids in maintaining the structural integrity from the input videos to the output videos. However, it’s observed that structural integrity alone doesn’t ensure motion accuracy, and SMA plays a crucial role in accurately capturing motion details.
Automatic Metrics User Study
Method Text-Align Temp-Con Edit-Acc Temp-Con Motion-Acc
Tune-A-Video 0.8289 0.8951 2.71 2.11 2.27
Tune-A-Video w/ Ours 0.8633 0.9568 3.41 2.88 3.25
ControlVideo 0.8686 0.9451 2.88 2.25 2.69
ControlVideo w/ Ours 0.8781 0.9590 3.56 3.16 3.40
Table 2: Quantitative evaluation of SMA within text-to-image based frameworks.
5.4 Quantitative Comparison.
Quantitative results are detailed in Tab. 2, following the metrics introduced in Sec. 4.2.4. Across both the Tune-A-Video and ControlVideo frameworks, the integration of SMA results in enhanced performance across all five evaluated metrics, notably achieving a substantial advantage in motion accuracy.
Figure 6: Visualization of (a) spatial frequency spectrum and (b) motion vectors estimated from the pre-trained Show-1 [45] without fine-tuning. (c) Ablation study on spectral motion alignment based on VMC [16].
6 Discussion
We explore the impact of motion vector refinement by examining motion vectors ($\delta{\boldsymbol{v}}_{0}$, $\delta\hat{{\boldsymbol{v}}}_{0}(t)$) and their (amplitude, phase) spectrum in Figure 6 ($t=700$). Figure 6b indicates that the original motion vector $\delta{\boldsymbol{v}}_{0}$ filters out appearance (e.g. background, color, etc) to some extent. However, it still faces frame-wise distortions or inconsistencies. For instance, in a specific example of this skateboarding man, non-motion-related artifacts, e.g. stair and fence patterns, background texture, etc., persist as distortions. These are characterized as high-frequency noises in the amplitude spectrum, Fig. 6a. Conversely, $\delta\hat{{\boldsymbol{v}}}_{0}^{n}(t)$ shows a reduction in high-frequency distortions while preserving essential motion information. As shown in (14), $\delta{{\boldsymbol{v}}}_{t}^{n}$ can be acquired from $\delta\boldsymbol{v}_{0}^{n}$ through the diffusion kernel. In this forward process, high-frequency components are perturbed more rapidly [20]. Consequently, the core motion information is consistently preserved in $\delta\boldsymbol{v}_{t}^{n}$ from coarse ($t\uparrow$) to fine ($t\downarrow$) level. Motivated by these insights, we prioritize low spatial frequency components during motion distillation to avoid overfitting to non-motion-related high-frequency distortions.
This motion vector refinement improves the overall fidelity as shown in Figure 6c. The application of VMC for motion transfer, without refinement, often fails to maintain motion patterns, introducing repetitive fence-like background distortions. Conversely, motion refinement within the frequency domain significantly improves fidelity and background detail. Moreover, global motion alignment further improves motion transfer. Specifically, while local motion alignment is generally effective, it occasionally generates "reversed" motions, i.e. an astronaut skateboarding in an upward direction. Note that this example is particularly challenging as it shows very subtle positional changes (video in project page). This highlights the limitations of conventional frameworks in identifying accurate motion from a single motion vector frame $\delta{\boldsymbol{v}}_{0}^{n}$, Figure 6b. In contrast, the proposed global motion alignment effectively mitigates these challenges, ensuring accurate learning of motion patterns. Fig. 7 further demonstrates the effectiveness of global motion alignment. Thus, our approach harmonizes both local and global motion alignment and effectively distills motion patterns while robustly reducing high-frequency distortions.
Figure 7: Ablation study on $\ell_{\text{global}}$. Global motion alignment facilitates motion transfer.
7 Conclusion
In this paper, we propose Spectral Motion Alignment (SMA), a novel frequency-based distillation framework for motion transfer. We explore the limitations of conventional motion estimation methods: (a) lack of global motion understanding, (b) vulnerability to spatial artifacts. By utilizing Fourier and wavelet transforms, SMA captures comprehensive motion dynamics, highlighting the significance of global motion context. We proved SMA’s effectiveness and compatibility through extensive validation across different motion transfer scenarios.
Reproducibility & Ethics Statement. For reproducibility, the code and datasets will be released to the public through the project page. Video generative models may pose potential risks for misuse in creating harmful content, which necessitates political regulation.
Appendix 0.A Pseudo Training Algorithm
In our work, we adopt the notation and expressions mostly from [16] for the preliminaries section and pseudo-code, due to its relevance to our focus on denoised motion vector estimates. We interchangeably use $\hat{{\boldsymbol{v}}}_{0}^{1:N}(t)$ and $\hat{{\boldsymbol{v}}}_{0}(t)$ in the main paper and Algorithm 1. While Algorithm 1 generalizes parameter $\theta$, each video diffusion model incorporates specific parameters such as $\theta_{\text{TA}}$ [16] and $\theta_{\text{LoRA}}$ [49].
Algorithm 1 Spectral Motion Alignment (SMA)
1:Input: $N$-frame input video sequence $({\boldsymbol{v}}_{0}^{n})_{n\in\{1,\dots,N\}}$, training prompt ${\mathcal{P}}$, textual encoder $\psi$, Training iterations $M$, Video diffusion models parameterized by $\theta$.
2:Output: Fine-tuned video diffusion models $\theta^{*}$.
3:
4:for $step=1$ to $M$ do
5:     Sample timestep $t\in[0,T]$ and Gaussian noise ${\boldsymbol{\epsilon}}_{t}^{1:N}$, where ${\boldsymbol{\epsilon}}_{t}^{n}\in\mathbb{R}^{d}\sim{\mathcal{N}}(0,I)$
6:     Prepare text embeddings $c=\psi({\mathcal{P}})$
7:
8:     1. Denoised motion vector estimation
9:     ${\boldsymbol{v}}_{t}^{1:N}=\sqrt{\bar{\alpha}_{t}}{\boldsymbol{v}}_{0}^{1:N}+\sqrt{1-\bar{\alpha}_{t}}{\boldsymbol{\epsilon}}_{t}^{1:N}$.
10:     $\hat{{\boldsymbol{v}}}_{0}^{1:N}(t)=\frac{1}{\sqrt{{\bar{\alpha}}_{t}}}\big{(}{\boldsymbol{v}}_{t}^{1:N}-\sqrt{1-{\bar{\alpha}}_{t}}{\boldsymbol{\epsilon}}_{\theta}({\boldsymbol{v}}_{t}^{1:N},t,c)\big{)}$.
11:
12:     2. Global motion alignment
13:     Conduct 1D DWT for each $s$-th pixel in $\delta{\boldsymbol{v}}_{0},\delta\hat{{\boldsymbol{v}}}_{0}(t)$ with Haar wavelet.
14:     $\ell_{\text{global}}(\delta{\boldsymbol{v}}_{0},\delta\hat{{\boldsymbol{v}}}_{0}(t))=\mathbb{E}_{t,s,j,k}\Big{[}\|{\mathcal{W}}_{\delta{\boldsymbol{v}}_{0,s}}(j,k)-{\mathcal{W}}_{\delta\hat{{\boldsymbol{v}}}_{0,s}(t)}(j,k)\|_{1}\Big{]}$.
15:
16:     3. Local motion refinement
17:     Obtain amplitude and phase spectrum for $\delta{\boldsymbol{v}}_{0}^{n}$ as $|\mathcal{F}_{\delta{\boldsymbol{v}}_{0}^{n}}(a,b)|,\angle\mathcal{F}_{\delta{\boldsymbol{v}}_{0}^{n}}(a,b)$.
18:     $\ell_{\text{l}ocal}^{a}(\delta{\boldsymbol{v}}_{0}^{n},\delta\hat{{\boldsymbol{v}}}_{0}^{n}(t))=\mathbb{E}_{t,n,a,b}\Big{[}{\mathcal{W}}(a,b)*\||\mathcal{F}_{\delta{\boldsymbol{v}}_{0}^{n}}(a,b)|-|\mathcal{F}_{\delta\hat{{\boldsymbol{v}}}_{0}^{n}(t)}(a,b)|\|_{1}\Big{]}$.
19:     $\ell_{\text{l}ocal}^{p}(\delta{\boldsymbol{v}}_{0}^{n},\delta\hat{{\boldsymbol{v}}}_{0}^{n}(t))=\mathbb{E}_{t,n,a,b}\Big{[}{\mathcal{W}}(a,b)*\|\angle\mathcal{F}_{\delta{\boldsymbol{v}}_{0}^{n}}(a,b)-\angle\mathcal{F}_{\delta\hat{{\boldsymbol{v}}}_{0}^{n}(t)}(a,b)\|_{1}\Big{]}$.
20:
21:     4. Overall optimization
22:     $\theta^{*}=\arg\min_{\theta}\mathbb{E}_{t,n,{\boldsymbol{\epsilon}}_{t}^{n},{\boldsymbol{\epsilon}}_{t}^{n+1}}\Big{[}\ell_{SMA}(\delta{\boldsymbol{v}}_{0},\delta\hat{{\boldsymbol{v}}}_{0}(t)\big{)}\Big{]}$ ($\ell_{SMA}$: objective in eq(22)).
23:end for
Appendix 0.B Related Work
0.B.1 Diffusion-based Video Editing
There has been considerable progress in adapting the achievements of diffusion-based image editing for video generative tasks. Compared to text-conditioned image generation, creating videos based solely on text introduces the complex challenge of producing temporally consistent and natural motion. In the absence of publicly available text-to-video diffusion models, Tune-A-Video [38] was at the forefront of one-shot based video editing. It proposes to inflate image diffusion model to pseudo video diffusion model by appending temporal modules to image diffusion model [28] and reprogramming spatial self-attention to spatio-temporal self-attention. Following this adaptation, the attention modules’ query projection matrices are fine-tuned on the input video. To eliminate the need for customizing model weights for every new video, various zero-shot editing methods have been developed. One approach involves guiding the generation process with attention maps, such as the injection of self-attention maps obtained from previous frames [5] or input video [25]. Another prevalent method integrates explicit structural cues, like depth or edge maps, into the reverse diffusion process. For instance, ControlNet [46] has been adapted for the video domain, facilitating the creation of structurally consistent frames in video generation [18] and translation tasks [15, 47]. Furthermore, GLIGEN’s [21] adaptation to the video domain by Ground-A-Video [17] demonstrates the use of both spatially-continuous depth map and spatially-discrete bounding box conditions, achieving multi-attribute editing of videos in a zero-shot manner.
The advent of open-source text-to-video diffusion models [34, 33, 6, 35] has spurred research into separating, altering, and combining the appearance and motion elements of videos [49, 16, 36, 1, 41, 42, 27]. MotionDirector [49] and DreamVideo [36] have each suggested approaches for dividing fine-tuning processes into distinct learning phases for subject appearance and temporal motion, utilizing efficient fine-tuning methods. On the other hand, VMC [16] focuses on distilling the motion within a video by calculating the residual vectors between consecutive frames. In their work, they fine-tune temporal attention layers within cascaded video diffusion models to synchronize the ground-truth motion vector with the denoised motion vector estimate, successfully generating videos that replicate the motion pattern of an input video within diverse visual scenarios. Similarly, [42] introduces a space-time feature loss that constructs self-similarity matrices based on the differences in attention features between frames. This approach aims to minimize the discrepancy in self-similarity between the input and output videos.
0.B.2 Frequency-aware Visual Generation
Spectral analysis plays a pivotal role in the domain of visual understanding and generation, offering insights into the temporal-spatial structure of pixel-domain frames through frequency-domain signals. [23] introduced a hierarchical motion estimation algorithm employing complex discrete wavelet transforms, effectively utilizing phase differences among subbands to indicate local translations within video frames. [29] enhanced scalable video compression through motion-compensated wavelet transforms, integrating a continuous deformable mesh motion model to achieve superior compression efficiency and motion representation.
Furthermore, these spectral insights have been instrumental in refining algorithms and deepening architectural understanding, particularly within the contexts of U-Net and autoencoder frameworks. [43] provided a groundbreaking reinterpretation of deep learning architectures for image reconstruction, establishing a connection between deep learning and classical signal processing theories, including wavelets and compressed sensing. [44] developed a wavelet-based correction method, WCT2, to augment photorealism in style transfer, leveraging whitening and coloring transforms to preserve structural integrity and statistics within the VGG feature space.
In the realm of U-Net in diffusion models, [37] highlighted the rapid dominance of noise over high-frequency information in residual U-Net denoisers. Concurrently, [19] proposed the Hybrid Video Diffusion Model (HVDM), a novel architecture that captures spatiotemporal dependencies using a disentangled representation combining 2D projection and 3D convolutions with wavelet decomposition. While [19] aims to reconstruct input video with frequency matching loss, our approach does not aim to reconstruct input, focusing instead on learning motion dynamics through motion vectors, thereby distinguishing our method within the landscape of spectral analysis applications in motion estimation and transfer.
Appendix 0.C Experimental details
For spectral global motion alignment, we mainly use Haar wavelets with the number of levels $l=3$ for 8-frame videos and $l=4$ for 16-frame videos. We use DWT1DForward function from the PyTorch package [^1] https://pytorch-wavelets.readthedocs.io/en/latest/index.html. For spectral local motion refinement, we fix $\delta=0.05$ in frequency domain weighting $\omega(a,b)$. We set $\lambda_{g}=0.4$ and $\lambda_{l}=0.2$ for many cases, where we recommend to fine-tune $\lambda_{g}$ in a range of $[0.2,0.5]$, and $\lambda_{g}$ in a range of $[0.1,0.3]$. We follow other configurations, e.g. optimization algorithm, learning rate, training steps, etc., from the original motion transfer frameworks.
Appendix 0.D Additional Results
This section provides additional qualitative comparisons of SMA across different baseline approaches. In Figure 8, we compare the efficacy of MotionDirector [49] in transferring motion, both with SMA integrated and without. Figures 9 and 10 delve into VMC’s [16] capabilities in customizing motion, showcasing outcomes with SMA and without its application. Additionally, Figure 11 contrasts the ability of Tune-A-Video and ControlVideo [38] to replicate the original motion, examining the impact of incorporating SMA. All our qualitative results can be viewed on our project page in video format.
Figure 8: Additional comparison within MotionDirector framework.
Figure 9: Additional comparison within VMC framework, deployed on Show-1 Cascade.
Figure 10: Additional comparison within VMC method, implemented on Zeroscope T2V.
Figure 11: Additional comparison within Tune-A-Video and ControlVideo frameworks.

### v2

@@
-
+Spectral Motion Alignment for Video Motion Transfer using Diffusion Models
 Abstract
-The evolution of diffusion models has greatly impacted video generation and understanding.
-Particularly, text-to-video diffusion models (VDMs) have significantly facilitated the customization of input video with target appearance, motion, etc.
-Despite these advances, challenges persist in accurately distilling motion information from video frames.
-While existing works leverage the consecutive frame residual as the target motion vector [16], they inherently lack global motion context and are vulnerable to frame-wise distortions.
-To address this, here we present Spectral Motion Alignment (SMA), a novel framework that refines and aligns motion vectors using Fourier and wavelet transforms.
-SMA learns motion patterns by incorporating frequency-domain regularization, facilitating the learning of whole-frame global motion dynamics, and mitigating spatial artifacts.
+Diffusion models have significantly facilitated the customization of input video with target appearance while maintaining its motion patterns.
+To distill the motion information from video frames, existing works often estimate motion representations as frame difference or correlation in pixel-/feature-space.
+Despite its simplicity, these methods have unexplored limitations, including lack of understanding of global motion context, and the introduction of motion-independent spatial distortions.
+To address this, we present Spectral Motion Alignment (SMA), a novel framework that refines and aligns motion representations in the spectral domain.
+Specifically, SMA learns spectral motion representations, facilitating the learning of whole-frame global motion dynamics, and effectively mitigating motion-independent artifacts.
 Extensive experiments demonstrate SMA’s efficacy in improving motion transfer while maintaining computational efficiency and compatibility across various video customization frameworks.
-Keywords:
-Diffusion models Video Motion transfer Wavelet transform
-Figure 1: One-shot Video Motion Transfer via Spectral Motion Alignment using Cascaded Video Diffusion.
-SMA facilitates the capture of long-range (top) and complex (bottom) motion patterns within videos.
-Visit here for comprehensive view of the videos.
+Figure 1: One-shot Video Motion Transfer via Spectral Motion Alignment using Cascaded Video Diffusion Models.
+SMA facilitates the capture of long-range (left) and complex (right) motion patterns within videos.
+Visit https://geonyeong-park.github.io/spectral-motion-alignment/ for a comprehensive view of the videos.
 1 Introduction
-The progression of diffusion models [31, 12] has played a pivotal role in advancing video generation and comprehension.
-Notably, recent advancements in text-conditioned Video Diffusion Models (VDMs, [13, 11, 30, 4, 9, 45, 3, 22]) have significantly enhanced the generation of highly realistic and superior-quality videos.
-Given the multifaceted nature of the video, encompassing motion dynamics, foreground object appearance, etc., several studies [49, 39, 16, 42, 41, 27, 36] aim to disentangle, modify, and compose these signals according to user intent.
-Additionally, various approaches specifically target to customize the motion, pose, shape, or appearance of real-world videos.
-In the context of motion customization, multiple studies explore how to distill motion information from pre-trained video diffusion models.
-Recent research suggests that motion patterns are inherently encoded in the underlying dependencies between latent frames or epsilon noises.
-As noted by [49], videos with similar motion tend to exhibit similar connectivity between latent frames.
-Additionally, [16] utilizes residual vectors between consecutive latent frames as "motion vectors," in line with optical flow principles, where residuals in consecutive frames represent motion dynamics.
-Specifically, they finetune temporal attention layers of pretrained VDM to align the ground-truth image-space residuals with their predicted denoised estimates.
-Their work reveals that aligning these predicted and ground-truth motion vectors corresponds to aligning epsilon (noise) residuals, readily obtainable from any off-the-shelf video diffusion models.
+Given the multifaceted nature of the video, encompassing motion dynamics, appearance, etc., several studies aim to disentangle and control these signals according to user intent.
+Recently, diffusion models (Sohl-Dickstein et al.
+2015; Ho, Jain, and Abbeel 2020) has played a pivotal role in video customization, owing to their superior sampling ability.
+In the context of motion customization using diffusion models, our goal is to transfer the motion patterns from an input video to the customized output video.
+This necessitates the accurate estimation and extraction of motion information from the input video.
+While fundamental techniques such as optical flow are effective for motion estimation, integrating these into diffusion models for customization is nontrivial.
+To address these challenges, recent researches suggest that motion patterns are inherently encoded in the underlying dependencies between frames or epsilon noises.
+For example, (Zhao et al.
+2023b) have observed that videos with similar motion tend to exhibit similar connectivity between latent frames.
+Additionally, (Jeong, Park, and Ye 2023) utilizes residual vectors between consecutive frames as "motion vectors," in line with optical flow principles, assuming frame residuals represent motion dynamics.
+Specifically, they finetune pretrained VDM to align the ground-truth pixel-space residuals with their predicted denoised estimates.
+Thus, these works leverage the pixel-space differences between input frames as a proxy of motion reference.
+While these motion representations can be obtained from off-the-shelf video diffusion models efficiently, current simple approximations have several adverse impacts.
+First, they may fail to capture the global context of motion.
+Since frame residuals may capture local motion patterns but are blind to whole-frame motion dynamics, for better motion dynamics modeling, we have to understand the whole-frame global context information during motion distillation.
+Furthermore, while the pixel- or feature-space residuals contain motion information, they may also contain inevitable disruptive variations that are unrelated to motion.
+These variations may include abrupt changes in the background, lightning, or other frame inconsistencies, leading to less reliable representation.
+To address these challenges, we introduce Spectral Motion Alignment (SMA), a novel framework for refining and aligning motion representations in the spectral domain, based on intuition that the motion may be well represented by its inherent frequency components.
+This framework includes two primary components: First, to capture the global motion context, we propose a spectral alignment loss between predicted and ground-truth motion vectors within the wavelet domain.
+This facilitates the learning of multi-scale motion dynamics by leveraging rich wavelet-domain representations of video considering the global frame transitions.
+Moreover, to mitigate the spatial artifacts and inconsistency in motion vectors, we propose 2D FFT-based motion vector refinement that aligns the amplitude and phase spectrum of ground truth and predicted motion vectors with prioritizing low-frequency components.
+This is because the high-frequency components in motion representations may be associated with frame-wise motion-independent artifacts (Figure 6).
+In summary, we encourage accurate motion transfer via harmonized global and local levels of spectral domain alignment.
+Our contributions are summarized as follows:
 Figure 2: Overview.
@@
 Our regularization includes (1) global motion alignment based on 1D wavelet-transform, and (2) local motion refinement based on 2D Fourier transform.
-While these efforts have made technical advances in motion distillation, the pitfalls of using latent frame residuals as motion vectors have not been fully studied.
-This paper focuses on the potential limitations of current motion estimation in diffusion models: (a) lack of global context information, and (b) vulnerability to frame-wise distortions.
-Since frame residuals may capture local motion patterns but are blind to whole-frame motion dynamics, for better motion dynamics modeling, we have to understand the whole-frame global context information during motion distillation.
-Furthermore, while the frame residuals contain motion information, they may also contain inevitable disruptive variations that are unrelated to motion.
-These variations may include abrupt changes in the background, distortions, consecutive frame inconsistencies, or lighting condition changes.
-To address these challenges, here we present Spectral Motion Alignment (SMA), a novel Fourier and wavelet-transform-based motion vector refinement and alignment framework.
-This includes two primary components: First, to learn the global motion context, we propose the spectral alignment loss between predicted and ground-truth motion vectors within the wavelet domain.
-This facilitates the learning of multi-scale motion dynamics by leveraging rich wavelet-domain representations of video considering the global frame transitions.
-Moreover, to mitigate the spatial artifacts and inconsistency in motion vectors, we propose 2D FFT-based motion vector refinement that aligns the amplitude and phase spectrum of ground truth and predicted motion vectors with prioritizing low-frequency components.
-This is because the high-frequency components in motion vectors may be associated with frame-wise non-motion-related artifacts (Figure 6).
-To sum up, we encourage accurate motion transfer via harmonized global and local levels of frequency-domain regularization.
-Our contributions are summarized as follows:
   • We introduce the Spectral Motion Alignment (SMA), a frequency-domain motion alignment framework that learns the underlying motion dynamics of input video via frequency-based regularization.
-Since conventional methods mostly rely on pixel-domain motion distillation, the proposed framework is orthogonal to existing motion customization models.
-  • SMA imposes negligible memory and computational burdens, as most off-the-shelf VDMs can readily compute estimates of motion vectors.
-For instance, VMC [16] with SMA demonstrates lightweight (15GB vRAM) and rapid (< 5 minutes) training.
-  • We validate the advantage of SMA across diverse motion patterns and subjects, and across various video motion transfer frameworks such as Video Diffusion-based [49], Cascaded Video Diffusion-based [16], Text-to-Image Diffusion-based [38], and ControlNet-based models [7].
+SMA is orthogonal and compatible to most of existing motion customization models as they often only rely on either pixel or feature space representations.
+  • SMA imposes negligible memory and computational burdens, as most off-the-shelf VDMs can readily compute motion vectors estimates.
+For instance, VMC (Jeong, Park, and Ye 2023) with SMA demonstrates lightweight (15GB vRAM) and rapid ($<$ 5 min) training.
+  • We validate the efficacy of SMA across diverse motion patterns, subjects, and various video motion transfer frameworks including Video Diffusion-based (Zhao et al.
+2023b), Cascaded Video Diffusion-based (Jeong, Park, and Ye 2023), T2I Diffusion-based (Wu et al.
+2023), and ControlNet-based models (Chen et al.
+2023b).
 2 Preliminaries
 2.1 Diffusion Models.
-Diffusion models [31, 12] generate samples from the Gaussian noise through reverse denoising processes.
-Given a clean sample ${\boldsymbol{x}}_{0}\sim p_{\operatorname{data}}({\boldsymbol{x}})$, the forward process is defined as a Markov chain with forward conditional densities
-$\begin{split}p({\boldsymbol{x}}_{t}\>|\>{\boldsymbol{x}}_{t-1})={\mathcal{N}}({\boldsymbol{x}}_{t}\>|\>\beta_{t}{\boldsymbol{x}}_{t-1},(1-\beta_{t})I)\\ p_{t}({\boldsymbol{x}}_{t}\>|\>{\boldsymbol{x}}_{0})={\mathcal{N}}({\boldsymbol{x}}_{t}\>|\>\sqrt{{\bar{\alpha}}}{\boldsymbol{x}}_{0},(1-{\bar{\alpha}})I),\end{split}$ (1)
-where ${\boldsymbol{x}}_{t}\in\mathbb{R}^{d}$ is a noisy latent variable at time $t$ that has the same dimension as ${\boldsymbol{x}}_{0}$, and $\beta_{t}$ denotes an increasing sequence of noise schedule where $\alpha_{t}\coloneqq 1-\beta_{t}$ and ${\bar{\alpha}}_{t}\coloneqq\Pi_{i=1}^{t}\alpha_{i}$.
-Then, the goal of diffusion model training is to obtain a residual denoiser ${\boldsymbol{\epsilon}}_{\theta^{*}}$:
-$\theta^{*}:=\mathop{\mathrm{argmin}}_{\theta}{\mathbb{E}}_{{\boldsymbol{x}}_{t}\sim p_{t}({\boldsymbol{x}}_{t}\>|\>{\boldsymbol{x}}_{0}),{\boldsymbol{x}}_{0}\sim p_{\operatorname{data}}({\boldsymbol{x}}_{0}),{\boldsymbol{\epsilon}}\sim{\mathcal{N}}(0,I)}\big{[}\left\lVert{\boldsymbol{\epsilon}}_{\theta}({\boldsymbol{x}}_{t},t)-{\boldsymbol{\epsilon}}\right\rVert\big{]}.$ (2)
+Diffusion models (Sohl-Dickstein et al.
+2015; Ho, Jain, and Abbeel 2020) generate samples from the Gaussian noise through reverse denoising processes.
+We denote a clean sample ${\boldsymbol{x}}_{0}\sim p_{\operatorname{data}}({\boldsymbol{x}})$, a noisy latent ${\boldsymbol{x}}_{t}\in\mathbb{R}^{d}$ at time $t$, $\beta_{t}$ as an increasing sequence of noise schedule, $\alpha_{t}\coloneqq 1-\beta_{t}$, and ${\bar{\alpha}}_{t}\coloneqq\Pi_{i=1}^{t}\alpha_{i}$.
+Then the goal of diffusion model training is to optimize a denoiser ${\boldsymbol{\epsilon}}_{\theta^{*}}$:
+$\theta^{*}:=\mathop{\mathrm{argmin}}_{\theta}{\mathbb{E}}_{{\boldsymbol{x}}_{t},{\boldsymbol{x}}_{0},{\boldsymbol{\epsilon}}}\big{[}\left\lVert{\boldsymbol{\epsilon}}_{\theta}({\boldsymbol{x}}_{t},t)-{\boldsymbol{\epsilon}}\right\rVert\big{]}.$ (1)
 The reverse sampling from $q({\boldsymbol{x}}_{t-1}|{\boldsymbol{x}}_{t},{\boldsymbol{\epsilon}}_{\theta^{*}}({\boldsymbol{x}}_{t},t))$ is then achieved by
@@
 where ${\boldsymbol{\epsilon}}\sim{\mathcal{N}}(0,{\boldsymbol{I}})$ and $\tilde{\beta}_{t}\coloneqq\frac{1-{\bar{\alpha}}_{t-1}}{1-{\bar{\alpha}}_{t}}\beta_{t}$.
-To accelerate sampling, DDIM [32] further proposes another sampling method as follows:
-${\boldsymbol{x}}_{t-1}=\sqrt{{\bar{\alpha}}_{t-1}}\hat{{\boldsymbol{x}}}_{0}(t)+\sqrt{1-{\bar{\alpha}}_{t-1}-\eta^{2}\tilde{\beta_{t}}^{2}}{\boldsymbol{\epsilon}}_{\theta^{*}}({\boldsymbol{x}}_{t},t)+\eta\tilde{\beta}_{t}{\boldsymbol{\epsilon}},$ (4)
-where $\eta\in[0,1]$ is a stochasticity parameter, and $\hat{{\boldsymbol{x}}}_{0}(t)$ is the denoised estimate which can be equivalently derived using Tweedie’s formula [8]:
+To accelerate sampling, DDIM (Song, Meng, and Ermon 2020) further proposes another sampling method as follows:
+${\boldsymbol{x}}_{t-1}=\sqrt{{\bar{\alpha}}_{t-1}}\hat{{\boldsymbol{x}}}_{0}(t)+\sqrt{1-{\bar{\alpha}}_{t-1}-\eta^{2}\tilde{\beta_{t}}^{2}}{\boldsymbol{\epsilon}}_{\theta^{*}}({\boldsymbol{x}}_{t},t)+\eta\tilde{\beta}_{t}{\boldsymbol{\epsilon}},$
+where $\eta\in[0,1]$ controls stochasticity, and $\hat{{\boldsymbol{x}}}_{0}(t)$ is the denoised estimate which can be equivalently derived using Tweedie’s formula (Efron 2011):
 $\hat{{\boldsymbol{x}}}_{0}(t)\coloneqq\frac{1}{\sqrt{{\bar{\alpha}}_{t}}}({\boldsymbol{x}}_{t}-\sqrt{1-{\bar{\alpha}}_{t}}{\boldsymbol{\epsilon}}_{\theta^{*}}({\boldsymbol{x}}_{t},t)).$ (3)
@@
 Throughout this paper, we will often omit $c$ from ${\boldsymbol{\epsilon}}_{\theta}({\boldsymbol{x}}_{t},t,c)$ if it does not lead to notational ambiguity.
-2.1.1 Video Diffusion Models.
-Video diffusion models [13, 11, 45] further attempt to model the video data distribution.
+Video Diffusion Models.
+Video diffusion models (Ho et al.
+2022b, a; Zhang et al.
+2023a) further attempt to model the video data distribution.
 Specifically, Let $({\boldsymbol{v}}^{n})_{n\in\{1,\dots,N\}}$ represents the $N$-frame input video sequence.
@@
 Spectral analysis techniques transform time-domain or pixel-domain signals (such as video frames) into the frequency domain, revealing the frequency components and their intensities.
-This transformation is valuable in video understanding as it helps in identifying repetitive motion patterns and underlying structures that may not be visible in the time or pixel domain.
 Fourier Transform. Let ${\boldsymbol{v}}^{n}\in\mathbb{R}^{H\times W}$ represents the $n$-th 2D video frame. Then, its frequency spectrum at coordinate $(a,b)$ is given as follows:
@@
 Then, the amplitude and phase is derived as follows:
-$|\mathcal{F}_{{\boldsymbol{v}}^{n}}(a,b)|=\sqrt{R(a,b)^{2}+I(a,b)^{2}},\angle\mathcal{F}_{{\boldsymbol{v}}^{n}}(a,b)=\arctan\Big{(}\frac{I(a,b)}{R(a,b)}\Big{)}.$ (8)
-Here, amplitude often indicates the existence of corresponding spatial frequencies within the image, often characterized by edges, textures, patterns, etc.
-Phase often encodes the structural spatial relationship.
+$\begin{split}|\mathcal{F}_{{\boldsymbol{v}}^{n}}(a,b)|=\sqrt{R(a,b)^{2}+I(a,b)^{2}},\\ \angle\mathcal{F}_{{\boldsymbol{v}}^{n}}(a,b)=\arctan\Big{(}\frac{I(a,b)}{R(a,b)}\Big{)}.\end{split}$ (6)
 Wavelet Transform.
-We first provide a brief introduction to the frame theory.
-Frames in signal processing are an extension of basis sets but with a relaxation that allows for redundancy.
-Let $\Psi=[\psi_{1}\dots\psi_{m}]\in\mathbb{R}^{n\times m}$, where $\{\psi_{k}\}_{k=1}^{m}$ is a set of functions in a Hilbert space $H$.
-Then, $\{\psi_{k}\}_{k=1}^{m}$ is called as a frame if it satisfies the following inequality:
-$\alpha\|f\|^{2}\leq\|\langle f,\Psi\rangle\|^{2}\leq\beta\|f\|^{2},\forall f\in H$ (9)
-where $c_{k}\coloneqq\langle f,\psi_{k}\rangle$ is a $k$-th frame expansion coefficient, and $\alpha,\beta$ are frame bounds.
-The original signal $f$ can be exactly reconstructed from expansion coefficients using the dual frame operator $\tilde{\Psi}$ which satisfies the so-called frame condition: $\tilde{\Psi}\Psi^{T}=I$.
 Wavelet frames, renowned for capturing multi-resolution scale features, are among the most prevalently utilized frame representations in signal processing.
@@
 $\mathcal{W}_{{\boldsymbol{v}}}(j,k)=\langle{\boldsymbol{v}}(t),\psi_{j,k}(t)\rangle.$ (9)
-Then, the original signal can be recovered from inverse DWT. In practice, this discrete wavelet transform can be implemented by convolution using an appropriate choice of filter bank.
+The original signal can be recovered from inverse DWT. In practice, this discrete wavelet transform can be implemented by convolution using an appropriate choice of filter bank.
 3 Spectral Motion Alignment
-Given an input video, our main goal is to develop a novel frequency-domain motion alignment framework that (a) distills the motion patterns $M^{*}$ of input video, and (b) transfers the motion patterns $M^{*}$ to output video within varied contexts, e.g.
-Cars with motion $M^{*}$rightarrow Tanks with motion $M^{*}$.
-Conventional motion distillation methods mostly rely on pixel-domain motion distillation objectives.
-However, frequency-domain analysis can further capture underlying motion patterns across a spectrum of frequency levels that mainly constitute motion.
-Accordingly, we propose Spectral Motion Alignment (SMA), a novel Fourier and wavelet-transform-based motion vector refinement and alignment framework.
-More details follow.
+Our main goal is to develop a novel spectral domain motion alignment framework that capture underlying complex motion patterns across a spectrum of frequency levels that mainly constitute motion.
+This is valuable in video understanding and customization as it helps in identifying repetitive motion patterns and underlying structures that may not be visible in the time or pixel domain.
+It is in orthogonal (and compatible) with conventional methods based on pixel- or feature-domain motion representations.
 3.1 Denoised Motion Vector Estimation
-To distill the motion information, we first estimate the motion vector of the input video following [16].
+To distill the motion information, we first estimate the initial motion representations (Jeong, Park, and Ye 2023; Zhao et al.
+2023b) in pixel space.
+For this, we follow VMC (Jeong, Park, and Ye 2023) as an representative example.
 The intuition is that residual vectors between consecutive frames may include information about the motion trajectories.
@@
 $\delta{\boldsymbol{v}}_{t}^{n}\coloneqq{\boldsymbol{v}}_{t}^{n+1}-{\boldsymbol{v}}_{t}^{n},$ (10)
-where the epsilon residual vector $\delta{\boldsymbol{\epsilon}}_{t}^{n}$ is similarly defined. This $\delta{\boldsymbol{v}}_{t}^{n}$ can be acquired through the following diffusion kernel [16]:
+where the epsilon residual vector $\delta{\boldsymbol{\epsilon}}_{t}^{n}$ is similarly defined. This $\delta{\boldsymbol{v}}_{t}^{n}$ can be acquired through the following diffusion kernel:
 $p(\delta{\boldsymbol{v}}_{t}^{n}\>|\>\delta{\boldsymbol{v}}_{0}^{n})={\mathcal{N}}(\delta{\boldsymbol{v}}_{t}^{n}\>|\>\sqrt{{\bar{\alpha}}_{t}}\delta{\boldsymbol{v}}_{0}^{n},2(1-{\bar{\alpha}}_{t})I).$ (11)
-Moreover, [16] shows that the ground-truth motion vector in pixel space $\delta{\boldsymbol{v}}_{0}^{n}$ can be derived as follows:
+Given that, the ground-truth motion vector in pixel space $\delta{\boldsymbol{v}}_{0}^{n}$ can be derived as follows:
 $\delta{\boldsymbol{v}}_{0}^{n}=\frac{1}{\sqrt{{\bar{\alpha}}_{t}}}\Big{(}\delta{\boldsymbol{v}}_{t}^{n}-\sqrt{1-{\bar{\alpha}}_{t}}\delta{\boldsymbol{\epsilon}}_{t}^{n}\Big{)}.$ (12)
-Similarly, one can obtain the denoised motion vector estimate $\delta\hat{{\boldsymbol{v}}}_{0}^{n}$ by using Tweedie’s formula as follows:
+Similarly, one can obtain the denoised estimate version of these motion representations $\delta\hat{{\boldsymbol{v}}}_{0}^{n}$ by using Tweedie’s formula as follows:
 $\hat{{\boldsymbol{v}}}_{0}^{1:N}(t)\coloneqq\frac{1}{\sqrt{{\bar{\alpha}}_{t}}}\big{(}{\boldsymbol{v}}_{t}^{1:N}-\sqrt{1-{\bar{\alpha}}_{t}}{\boldsymbol{\epsilon}}_{\theta}({\boldsymbol{v}}_{t}^{1:N},t)\big{)},$ (13)
 where $\hat{{\boldsymbol{v}}}_{0}^{1:N}(t)$ is an empirical Bayes optimal posterior expectation $\mathbb{E}[{\boldsymbol{v}}_{0}^{1:N}\>|\>{\boldsymbol{v}}_{t}^{1:N}]$.
-In the context of motion transfer, the authors in [16] finetune the temporal attention layers of VDM by aligning each motion vector $\delta{\boldsymbol{v}}_{0}^{n}$ and its denoised estimate $\delta\hat{{\boldsymbol{v}}}_{0}^{n}(t)$:
+In the context of motion transfer, we aim to align the ground-truth and estimated motion vectors by fine-tuning the pre-trained VDM (Jeong, Park, and Ye 2023; Zhao et al. 2023b):
 $\min_{\theta}\mathbb{E}_{t,n,{\boldsymbol{\epsilon}}^{t,n},{\boldsymbol{\epsilon}}^{t,n+1}}\Big{[}\ell_{\text{align}}\big{(}\delta{\boldsymbol{v}}_{0}^{n},\delta\hat{{\boldsymbol{v}}}_{0}^{n}(t)\big{)}\Big{]}.$ (14)
-While these advancements in motion distillation mark significant progress, Figure 1, 4 and 6 indicate that [16] still has potential for further refinement.
+While these advancements in motion distillation mark significant progress, Figure 1, 4 and 6 indicate that existing methods still has potential for further refinement.
 3.2 Spectral Global Motion Alignment
 One of the primary limitations in (14) is that it may not fully encapsulate the global motion dynamics.
-This is partly because [16] employs ‘frame-wise’ matching losses, such as $\ell_{2}$-distance.
-While this is effective, it may overlook the comprehensive motion dynamics by focusing solely on pairwise frame comparisons.
+Specifically, it locally focuses on pairwise frame comparisons which may lead to overlooking the comprehensive motion dynamics of given object overall frames.
 To mitigate these problems, we explore the use of wavelet transforms in motion distillation. In this paper, we use Haar wavelet, whose low and high pass filters are given as follows:
@@
 Then, the frequency-matching loss between $\delta{\boldsymbol{v}}_{0}$ and $\delta\hat{{\boldsymbol{v}}}_{0}(t)$ is defined with DWT in (9) as follows:
-$\ell_{\text{global}}(\delta{\boldsymbol{v}}_{0},\delta\hat{{\boldsymbol{v}}}_{0}(t))=\mathbb{E}_{t,s,j,k}\Big{[}\|{\mathcal{W}}_{\delta{\boldsymbol{v}}_{0,s}}(j,k)-{\mathcal{W}}_{\delta\hat{{\boldsymbol{v}}}_{0,s}(t)}(j,k)\|_{1}\Big{]}.$ (19)
+$\begin{split}\ell_{\text{global}}&(\delta{\boldsymbol{v}}_{0},\delta\hat{{\boldsymbol{v}}}_{0}(t))=\\ &\mathbb{E}_{t,s,j,k}\Big{[}\|{\mathcal{W}}_{\delta{\boldsymbol{v}}_{0,s}}(j,k)-{\mathcal{W}}_{\delta\hat{{\boldsymbol{v}}}_{0,s}(t)}(j,k)\|_{1}\Big{]}.\end{split}$ (16)
 Considering that the wavelet transform allows multi-resolution analysis of motion vectors, it enables us to handle motions at various scales and frequencies effectively.
@@
 3.3 Spectral Local Motion Refinement
-Another problem in (17) is that the residuals between consecutive frames, while rich in motion information, may also encapsulate high-frequency local distortions, background noise, and other non-motion-related artifacts.
-Accordingly, we focus on particularly prioritizing low-to-moderate spatial frequency components.
-Specifically, following the amplitude and phase spectrum definition in (8), we define amplitude and phase matching loss as follows:
-$\begin{split}&\ell_{\text{l}ocal}^{A}(\delta{\boldsymbol{v}}_{0}^{n},\delta\hat{{\boldsymbol{v}}}_{0}^{n}(t))=\mathbb{E}_{t,n,a,b}\Big{[}\omega(a,b)\||\mathcal{F}_{\delta{\boldsymbol{v}}_{0}^{n}}(a,b)|-|\mathcal{F}_{\delta\hat{{\boldsymbol{v}}}_{0}^{n}(t)}(a,b)|\|_{1}\Big{]},\\ &\ell_{\text{l}ocal}^{P}(\delta{\boldsymbol{v}}_{0}^{n},\delta\hat{{\boldsymbol{v}}}_{0}^{n}(t))=\mathbb{E}_{t,n,a,b}\Big{[}\omega(a,b)\|\angle\mathcal{F}_{\delta{\boldsymbol{v}}_{0}^{n}}(a,b)-\angle\mathcal{F}_{\delta\hat{{\boldsymbol{v}}}_{0}^{n}(t)}(a,b)\|_{1}\Big{]},\end{split}$ (20)
+Another problem in (14) is that the estimated motion representations may encapsulate high-frequency local distortions, background noise, and other non-motion-related artifacts.
+By aligning the denoised estimates with these artifacts, the fine-tuned VDM may erroneously reproduce similar high-frequency artifacts as in Figure 6.
+Accordingly, we focus on prioritizing low-to-moderate spatial frequency components particularly.
+Specifically, following the amplitude and phase spectrum definition in (6), we define amplitude and phase matching loss, $\ell_{\text{l}ocal}^{A}(\delta{\boldsymbol{v}}_{0}^{n},\delta\hat{{\boldsymbol{v}}}_{0}^{n}(t))$ and $\ell_{\text{l}ocal}^{P}(\delta{\boldsymbol{v}}_{0}^{n},\delta\hat{{\boldsymbol{v}}}_{0}^{n}(t))$, as follows:
+$\begin{split}&(A)\ \mathbb{E}_{t,n,a,b}\Big{[}\omega(a,b)\||\mathcal{F}_{\delta{\boldsymbol{v}}_{0}^{n}}(a,b)|-|\mathcal{F}_{\delta\hat{{\boldsymbol{v}}}_{0}^{n}(t)}(a,b)|\|_{1}\Big{]},\\ &(P)\ \mathbb{E}_{t,n,a,b}\Big{[}\omega(a,b)\|\angle\mathcal{F}_{\delta{\boldsymbol{v}}_{0}^{n}}(a,b)-\angle\mathcal{F}_{\delta\hat{{\boldsymbol{v}}}_{0}^{n}(t)}(a,b)\|_{1}\Big{]},\end{split}$ (17)
 where the frequency domain weighting $\omega(a,b)$ is defined as
-$\omega(a,b)=\Big{[}\big{(}\frac{H}{2}\big{)}^{2}+\big{(}\frac{W}{2}\big{)}^{2}\Big{]}^{\delta}-\Big{[}\big{(}a-\frac{H}{2}\big{)}^{2}+\big{(}b-\frac{W}{2}\big{)}^{2}\Big{]}^{\delta}+1$ (21)
-for $0<a<H,0<b<W$, and otherwise, set to zero. This introduces a weighting [40] that prioritizes low-frequency components for $\delta>0$.
+$\begin{split}\omega(a,b)=\Big{[}\big{(}\frac{H}{2}\big{)}^{2}+\big{(}\frac{W}{2}\big{)}^{2}\Big{]}^{\delta}-\Big{[}\big{(}a-\frac{H}{2}\big{)}^{2}+\big{(}b-\frac{W}{2}\big{)}^{2}\Big{]}^{\delta}+1\end{split}$
+for $0<a<H,0<b<W$, and otherwise, set to zero. This introduces a weighting (Yang et al. 2022) that prioritizes low-frequency components for $\delta>0$.
 Figure 3: Comparison within MotionDirector framework.
@@
 To sum up, the overall spectral motion alignment framework is given as follows:
-$\min_{\theta}\mathbb{E}_{t,n,{\boldsymbol{\epsilon}}_{t}^{n},{\boldsymbol{\epsilon}}_{t}^{n+1}}\Big{[}\ell_{\text{align}}\big{(}\delta{\boldsymbol{v}}_{0}^{n},\delta\hat{{\boldsymbol{v}}}_{0}^{n}(t)\big{)}+\lambda_{g}\ell_{\text{global}}\big{(}\delta{\boldsymbol{v}}_{0},\delta\hat{{\boldsymbol{v}}}_{0}(t)\big{)}+\lambda_{l}\ell_{\text{l}ocal}\big{(}\delta{\boldsymbol{v}}_{0}^{n},\delta\hat{{\boldsymbol{v}}}_{0}^{n}(t)\big{)}\Big{]},$ (22)
+$\begin{split}\min_{\theta}&\mathbb{E}_{t,n,{\boldsymbol{\epsilon}}_{t}^{n},{\boldsymbol{\epsilon}}_{t}^{n+1}}\Big{[}\ell_{\text{align}}\big{(}\delta{\boldsymbol{v}}_{0}^{n},\delta\hat{{\boldsymbol{v}}}_{0}^{n}(t)\big{)}+\\ &\lambda_{g}\ell_{\text{global}}\big{(}\delta{\boldsymbol{v}}_{0},\delta\hat{{\boldsymbol{v}}}_{0}(t)\big{)}+\lambda_{l}\ell_{\text{l}ocal}\big{(}\delta{\boldsymbol{v}}_{0}^{n},\delta\hat{{\boldsymbol{v}}}_{0}^{n}(t)\big{)}\Big{]},\end{split}$ (18)
 […]
@@
 While diverse diffusion-based motion distillation frameworks adopt their pixel-domain motion learning objectives, the proposed frequency-domain alignment seamlessly integrates with these arbitrary objectives.
-Moreover, different motion distillation frameworks target specific parameters $\theta$ for fine-tuning, varying from temporal attention layers [16] to dual-path LoRAs [49].
-We empirically demonstrate global compatibility of the proposed spectral motion alignment with diverse neural architectures and parameterizations.
+Moreover, different motion distillation frameworks target specific parameters $\theta$ for fine-tuning, varying from temporal attention layers (Jeong, Park, and Ye 2023) to dual-path LoRAs (Zhao et al.
+2023b).
+We empirically demonstrate the global compatibility of the proposed spectral motion alignment with diverse neural architectures and parameterizations.
 Pseudo-code is provided in the appendix.
-Figure 4: Comparison within VMC framework, using Show-1 cascaded model (top) and Zeroscope (bottom) non-cascaded model. Visit our project page for an in-depth view.
+3.5 Extending SMA to Feature Space
+Beyond pixel-space motion representations, SMA can be further extended to semantic diffusion features (DIFT, (Tang et al.
+2023)).
+Specifically, Diffusion-Motion-Transfer (DMT, (Yatim et al.
+2023)) constructs motion vectors based on pairwise differences in space-time diffusion features, which are then utilized for latent optimization-based video motion transfer.
+Given input and target video latents, ${\boldsymbol{v}}_{t}$ and $\tilde{{\boldsymbol{v}}}_{t}$, the model extracts space-time features $f({\boldsymbol{v}}_{t})$ and $f(\tilde{{\boldsymbol{v}}}_{t})$.
+Then, the feature residuals are defined as $\delta f({\boldsymbol{v}}_{t})^{n}$ and $\delta f(\tilde{{\boldsymbol{v}}}_{t})^{n}$, the $n$-th consecutive difference between hidden feature frames.
+This leads to the spectral alignment objective in feature space as follows:
+$\begin{split}\mathbb{E}\Big{[}&\ell_{\text{DMT}}\big{(}f({\boldsymbol{v}}_{t}),f(\tilde{{\boldsymbol{v}}}_{t})\big{)}+\\ &\lambda_{g}\ell_{\text{global}}\big{(}\delta f({\boldsymbol{v}}_{t}),\delta f(\tilde{{\boldsymbol{v}}}_{t})\big{)}+\lambda_{l}\ell_{\text{l}ocal}\big{(}\delta f({\boldsymbol{v}}_{t})^{n},\delta f(\tilde{{\boldsymbol{v}}}_{t})^{n}\big{)}\Big{]},\end{split}$ (19)
+where $\ell_{\text{DMT}}$ refers to the original space-time feature loss in (Yatim et al.
+2023).
+Note that DMT does not finetune the models, leveraging (19) for a latent optimization in sampling process.
+We demonstrate the effectiveness of spectral alignment in the diffusion feature space by comparing it against the original DMT framework in Fig.4-bottom.
+Figure 4: Comparison within VMC framework using Show-1 video model (top) and DMT framework using Zeroscope video model (bottom).
+Each demonstrate the compatibility of SMA in pixel-space and feature-space, respectively.
 4 Experiments using T2V Diffusion Models
 4.1 Experimental Setting
-To assess the capability of Spectral Motion Alignment (SMA) to capture accurate motion within contemporary diffusion-based motion learning frameworks, we curated a dataset comprising 30 text-video pairs sourced from the publicly available DAVIS [24] and WebVid-10M [2] collections.
+To assess the capability of Spectral Motion Alignment (SMA) to capture accurate motion within contemporary diffusion-based motion learning frameworks, we curated a dataset comprising 30 text-video pairs sourced from the publicly available DAVIS (Pont-Tuset et al.
+2017) and WebVid-10M (Bain et al.
+2021) collections.
 This dataset is deliberately designed to cover a broad spectrum of motion types and subjects, with video lengths ranging between 8 and 16 frames.
-For this study, we leverage two foundational text-to-video diffusion models: Zeroscope [33], a non-cascaded VDM, and Show-1 [45], a cascaded VDM.
+For this study, we leverage two foundational text-to-video diffusion models: Zeroscope (Sterling 2023), a non-cascaded VDM, and Show-1 (Zhang et al.
+2023a), a cascaded VDM.
 More details are provided in appendix.
 4.2 Baselines
-4.2.1 MotionDirector.
-MotionDirector [49] pioneered the concept of motion customization within video diffusion literature.
-To distinctively tailor the appearance and motion of a video, [49] developed a unique dual-path framework employing Low-Rank Adaptation (LoRA, [14]).
-Within this architecture, spatial LoRAs are optimized using the diffusion training loss on a single, randomly chosen frame, rather than on the full video sequence.
-Conversely, temporal LoRAs are trained by considering all frames of the video, using an anchor frame based appearance-debiased temporal loss.
-4.2.2 VMC.
-VMC [16] achieves state-of-the-art performance in motion customization through their novel epsilon residual matching objective, facilitating efficient motion distillation within cascaded video diffusion framework.
-In their work, they showcased the efficient of motion distillation within cascaded video diffusion models, by optimizing only keyframe generation model while freezing later modules.
-As VMC represents the most closely related work to ours, we conduct a two-fold validation within this baseline, including both cascaded and non-cascaded video diffusion scenarios.
-4.2.3 Qualitative Comparison.
+MotionDirector (Zhao et al.
+2023b) tailor the appearance and motion of a video by developing a unique dual-path (spatial, temporal) framework with Low-Rank Adaptation (LoRA, (Hu et al.
+2021)).
+VMC (Jeong, Park, and Ye 2023) achieves state-of-the-art performance in motion customization through their novel epsilon residual matching objective, facilitating efficient motion distillation within a cascaded video diffusion.
+DMT (Yatim et al.
+2023) proposes a new space-time feature loss, guiding the sampling process towards preserving the motion patterns while complying with the target object.
+Please see Sec 3.5 for more details.
+4.3 Qualitative Comparison.
 Fig.
-3 offers a qualitative analysis comparing MotionDirector’s performance with and without the implementation of SMA.
-Likewise, Fig.
-4 examines VMC’s performance with and without SMA implementation.
-Specifically, videos produced using a cascaded diffusion pipeline are displayed in Fig.
-4-top, whereas those generated through a non-cascaded diffusion model are shown in the bottom.
-Videos generated without SMA correctly capture the appearance as dictated by the target text, but fail to accurately mirror the original video’s motion patterns.
-In contrast, SMA significantly enhances motion understanding, precisely distinguishing between dynamic and static elements in the scene.
+3 and 4 offer qualitative comparisons with and without SMA.
+The top of Figure 4 shows videos from a cascaded diffusion pipeline, while the bottom displays those from a non-cascaded model.
+Without SMA, videos may capture appearance to some extent but fail to replicate motion patterns accurately.
+In contrast, SMA significantly improves motion transfer, distinguishing dynamic from static objects.
 For instance, in the last example of Fig.
-3, MotionDirector with SMA produces a video where only the eagle moves from right to left, whereas without SMA, the video inaccurately depicts the ground moving alongside the eagle.
-4.2.4 Quantitative Comparison.
+3, SMA produces a video where only the eagle moves from right to left, whereas without SMA, the video inaccurately depicts the ground moving alongside the eagle.
+4.4 Quantitative Comparison.
 The results of our quantitative evaluation are presented in Table 1.
-We employ CLIP encoders [26] for automated evaluation.
-To evaluate text-video alignment [10], we measure the average cosine similarity between the target text prompt and the frames generated.
+To evaluate text-video alignment (Hessel et al.
+2021), we measure the average cosine similarity between the target text prompt and the frames generated.
 Regarding frame consistency, we extract CLIP image features for each frame in the output video and subsequently calculate the average cosine similarity among all frame pairs in the video.
@@
 Table 1: Quantitative evaluation of SMA within text-to-video based frameworks.
-Figure 5: Comparison within Tune-A-Video and ControlVideo (Depth) baseline.
-5 Experiments using T2I Video Diffusion Models
+Figure 5: Comparison within Tune-A-Video (Top) and ControlVideo-Depth (Bottom) baseline.
+Figure 6: Visualization of (a) spatial frequency spectrum and (b) motion vectors estimated from the pre-trained Show-1 (Zhang et al.
+2023a) without fine-tuning.
+(c) Ablation study on spectral motion alignment based on VMC (Jeong, Park, and Ye 2023).
+5 T2I Video Diffusion Models
 5.1 Experimental Setting
-To assess the effectiveness of SMA within methodologies built on text-to-image diffusion model, we use the same text-video pairs previously assembled in Sec.
+We further evaluate the efficacy of SMA with methods based on text-to-image diffusion model.
+Same text-video pairs are used as in Sec.
 4.1.
 The resolution for all produced videos is standardized to 512x512.
-In this experiment, Stable Diffusion v1-5 [28] and ControlNet-Depth [46] are utilized.
+In this experiment, Stable Diffusion v1-5 (Rombach et al.
+2022) and ControlNet-Depth (Zhang, Rao, and Agrawala 2023) are utilized.
 5.2 Baselines
-5.2.1 Tune-A-Video.
-Tune-A-Video [38] was at the forefront of image-diffusion-based video editing at a time when video diffusion models were not yet publicly accessible.
-It begins with transforming a pretrained T2I diffusion model to psuedo T2V model by adding temporal attention layers and expanding spatial self-attention into spatio-temporal attention.
-Attention projection layers within this pseudo T2V model are then fine-tuned on a specific input video, utilizing the diffusion training objective (Eq.
-2).
-5.2.2 ControlVideo.
-ControlVideo [48] is another one-shot-based video editing method, starting from pretrained T2I model.
-ControlVideo extends ControlNet [46] from image to video to incorporate structural cues obtained from the input video.
-Subsequently, akin to Tune-A-Video [38], attention projection layers of the backbone U-Net and ControlNet are customized on the input video.
+Tune-A-Video (Wu et al.
+2023) transforms a pretrained T2I diffusion model to psuedo T2V model by adding temporal attention layers and expanding spatial self-attention into spatio-temporal attention.
+ControlVideo (Zhao et al.
+2023a) is another one-shot-based video editing method stems from pretrained T2I model.
+ControlVideo extends ControlNet (Zhang, Rao, and Agrawala 2023) from image to video to incorporate structural cues obtained from the input video.
 5.3 Qualitative Comparison.
 Fig.
-5-(top) demonstrates the qualitative benefits of incorporating SMA into the Tune-A-Video method.
-Without SMA, Tune-A-Video often experiences flickering in foreground objects, a problem that SMA significantly reduces.
+5-(top) demonstrates the efficacy of SMA with Tune-A-Video method, where SMA alleviates the flickering artifacts in foreground objects.
 Fig.
-5-(bottom) illustrates the qualitative improvements brought by integrating SMA with the ControlVideo framework.
-In ControlVideo, depth control aids in maintaining the structural integrity from the input videos to the output videos.
-However, it’s observed that structural integrity alone doesn’t ensure motion accuracy, and SMA plays a crucial role in accurately capturing motion details.
+5-(bottom) further illustrates the improvements with the ControlVideo framework.
+While depth control encourages ControlVideo to maintain the structural integrity, Fig.
+5 shows that it is not sufficient for motion accuracy, where SMA plays a crucial role in accurate capture of motion details.
 Automatic Metrics User Study
@@
 2, following the metrics introduced in Sec.
-4.2.4.
-Across both the Tune-A-Video and ControlVideo frameworks, the integration of SMA results in enhanced performance across all five evaluated metrics, notably achieving a substantial advantage in motion accuracy.
-Figure 6: Visualization of (a) spatial frequency spectrum and (b) motion vectors estimated from the pre-trained Show-1 [45] without fine-tuning.
-(c) Ablation study on spectral motion alignment based on VMC [16].
-6 Discussion
-We explore the impact of motion vector refinement by examining motion vectors ($\delta{\boldsymbol{v}}_{0}$, $\delta\hat{{\boldsymbol{v}}}_{0}(t)$) and their (amplitude, phase) spectrum in Figure 6 ($t=700$).
-Figure 6b indicates that the original motion vector $\delta{\boldsymbol{v}}_{0}$ filters out appearance (e.g.
-background, color, etc) to some extent.
-However, it still faces frame-wise distortions or inconsistencies.
-For instance, in a specific example of this skateboarding man, non-motion-related artifacts, e.g.
-stair and fence patterns, background texture, etc., persist as distortions.
+4.4.
+Across both the Tune-A-Video and ControlVideo frameworks, the integration of SMA improves performance across all five evaluated metrics, notably achieving a substantial advantage in motion accuracy.
+6 Analysis
+We explore the impact of SMA by examining motion vectors ($\delta{\boldsymbol{v}}_{0}$, $\delta\hat{{\boldsymbol{v}}}_{0}(t)$) and their (amplitude, phase) spectrum in Figure 6 ($t=700$).
+Figure 6b indicates that the pixel-space motion vector $\delta{\boldsymbol{v}}_{0}$ faces frame-wise distortions or inconsistencies.
+For instance, the motion-independent artifacts, e.g.
+stair and fence patterns, background texture, etc., persist as distortions in the obtained motion vectors.
 These are characterized as high-frequency noises in the amplitude spectrum, Fig.
 6a.
-Conversely, $\delta\hat{{\boldsymbol{v}}}_{0}^{n}(t)$ shows a reduction in high-frequency distortions while preserving essential motion information.
-As shown in (14), $\delta{{\boldsymbol{v}}}_{t}^{n}$ can be acquired from $\delta\boldsymbol{v}_{0}^{n}$ through the diffusion kernel.
-In this forward process, high-frequency components are perturbed more rapidly [20].
-Consequently, the core motion information is consistently preserved in $\delta\boldsymbol{v}_{t}^{n}$ from coarse ($t\uparrow$) to fine ($t\downarrow$) level.
-Motivated by these insights, we prioritize low spatial frequency components during motion distillation to avoid overfitting to non-motion-related high-frequency distortions.
-This motion vector refinement improves the overall fidelity as shown in Figure 6c.
-The application of VMC for motion transfer, without refinement, often fails to maintain motion patterns, introducing repetitive fence-like background distortions.
-Conversely, motion refinement within the frequency domain significantly improves fidelity and background detail.
+Motivated by these observations, we prioritize low spatial frequency components during motion distillation to avoid overfitting to motion-independent high-frequency distortions.
+This representations refinement improves the overall fidelity and removes background distortions (Figure 6c).
 Moreover, global motion alignment further improves motion transfer.
-Specifically, while local motion alignment is generally effective, it occasionally generates "reversed" motions, i.e.
+Specifically, without considering the global motion dynamics, existing frameworks occasionally generate “reversed" motions, i.e.
 an astronaut skateboarding in an upward direction.
-Note that this example is particularly challenging as it shows very subtle positional changes (video in project page).
-This highlights the limitations of conventional frameworks in identifying accurate motion from a single motion vector frame $\delta{\boldsymbol{v}}_{0}^{n}$, Figure 6b.
+This highlights the limitations of conventional frameworks in understanding accurate motion from a single frame of motion vector $\delta{\boldsymbol{v}}_{0}^{n}$.
 In contrast, the proposed global motion alignment effectively mitigates these challenges, ensuring accurate learning of motion patterns.
-Fig.
-7 further demonstrates the effectiveness of global motion alignment.
-Thus, our approach harmonizes both local and global motion alignment and effectively distills motion patterns while robustly reducing high-frequency distortions.
-Figure 7: Ablation study on $\ell_{\text{global}}$. Global motion alignment facilitates motion transfer.
+Table 3 demonstrates the importance of both global/local terms (More qualitative ablation results in appendix).
+Baseline Baseline $+\mathcal{L}_{\text{local}}$ Baseline $+\mathcal{L}_{\text{global}}$ Baseline $+\mathcal{L}_{\text{local}}+\mathcal{L}_{\text{global}}$
+Text-Align 0.8163 0.8382 0.8344 0.8427 (+ 0.026)
+Frame-Con 0.9497 0.9613 0.9585 0.9659 (+ 0.016)
+Table 3: Quantitative ablation of $\mathcal{L}_{\text{local}}$ and $\mathcal{L}_{\text{global}}$.
 7 Conclusion
-In this paper, we propose Spectral Motion Alignment (SMA), a novel frequency-based distillation framework for motion transfer.
+We propose Spectral Motion Alignment (SMA), a novel motion distillation framework in spectral domain.
 We explore the limitations of conventional motion estimation methods: (a) lack of global motion understanding, (b) vulnerability to spatial artifacts.
-By utilizing Fourier and wavelet transforms, SMA captures comprehensive motion dynamics, highlighting the significance of global motion context.
-We proved SMA’s effectiveness and compatibility through extensive validation across different motion transfer scenarios.
-Reproducibility & Ethics Statement.
-For reproducibility, the code and datasets will be released to the public through the project page.
-Video generative models may pose potential risks for misuse in creating harmful content, which necessitates political regulation.
-Appendix 0.A Pseudo Training Algorithm
-In our work, we adopt the notation and expressions mostly from [16] for the preliminaries section and pseudo-code, due to its relevance to our focus on denoised motion vector estimates.
+Then, we mitigate these problems by harmonizing both local and global motion alignment and effectively distills motion patterns.
+Acknowledgments
+This work was supported by Institute of Information &\&& communications Technology Planning &\&& Evaluation (IITP) grant funded by the Korea government(MSIT) (No.2019-0-00075, Artificial Intelligence Graduate School Program(KAIST), No.
+RS-2023-00233251, System3 reinforcement learning with high-level brain functions), National Research foundation of Korea(NRF) (**RS-2023-00262527**, RS-2024-00336454, RS-2024-00341805).
+Appendix A Pseudo Training Algorithm
+In our work, we adopt the notation and expressions mostly from (Jeong, Park, and Ye 2023) for the preliminaries section and pseudo-code, due to its relevance to our focus on denoised motion vector estimates.
 We interchangeably use $\hat{{\boldsymbol{v}}}_{0}^{1:N}(t)$ and $\hat{{\boldsymbol{v}}}_{0}(t)$ in the main paper and Algorithm 1.
-While Algorithm 1 generalizes parameter $\theta$, each video diffusion model incorporates specific parameters such as $\theta_{\text{TA}}$ [16] and $\theta_{\text{LoRA}}$ [49].
+While Algorithm 1 generalizes parameter $\theta$, each video diffusion model incorporates specific parameters such as $\theta_{\text{TA}}$ (Jeong, Park, and Ye 2023) and $\theta_{\text{LoRA}}$ (Zhao et al.
+2023b).
 Algorithm 1 Spectral Motion Alignment (SMA)
@@
 23:end for
-Appendix 0.B Related Work
-0.B.1 Diffusion-based Video Editing
+Appendix B Related Work
+B.1 Diffusion-based Video Editing
 There has been considerable progress in adapting the achievements of diffusion-based image editing for video generative tasks.
 Compared to text-conditioned image generation, creating videos based solely on text introduces the complex challenge of producing temporally consistent and natural motion.
-In the absence of publicly available text-to-video diffusion models, Tune-A-Video [38] was at the forefront of one-shot based video editing.
-It proposes to inflate image diffusion model to pseudo video diffusion model by appending temporal modules to image diffusion model [28] and reprogramming spatial self-attention to spatio-temporal self-attention.
+In the absence of publicly available text-to-video diffusion models, Tune-A-Video (Wu et al.
+2023) was at the forefront of one-shot based video editing.
+It proposes to inflate image diffusion model to pseudo video diffusion model by appending temporal modules to image diffusion model (Rombach et al.
+2022) and reprogramming spatial self-attention to spatio-temporal self-attention.
 Following this adaptation, the attention modules’ query projection matrices are fine-tuned on the input video.
 To eliminate the need for customizing model weights for every new video, various zero-shot editing methods have been developed.
-One approach involves guiding the generation process with attention maps, such as the injection of self-attention maps obtained from previous frames [5] or input video [25].
+One approach involves guiding the generation process with attention maps, such as the inekction of self-attention maps obtained from the input video (Ceylan, Huang, and Mitra 2023; Qi et al.
+2023).
 Another prevalent method integrates explicit structural cues, like depth or edge maps, into the reverse diffusion process.
-For instance, ControlNet [46] has been adapted for the video domain, facilitating the creation of structurally consistent frames in video generation [18] and translation tasks [15, 47].
-Furthermore, GLIGEN’s [21] adaptation to the video domain by Ground-A-Video [17] demonstrates the use of both spatially-continuous depth map and spatially-discrete bounding box conditions, achieving multi-attribute editing of videos in a zero-shot manner.
-The advent of open-source text-to-video diffusion models [34, 33, 6, 35] has spurred research into separating, altering, and combining the appearance and motion elements of videos [49, 16, 36, 1, 41, 42, 27].
-MotionDirector [49] and DreamVideo [36] have each suggested approaches for dividing fine-tuning processes into distinct learning phases for subject appearance and temporal motion, utilizing efficient fine-tuning methods.
-On the other hand, VMC [16] focuses on distilling the motion within a video by calculating the residual vectors between consecutive frames.
+For instance, ControlNet (Zhang, Rao, and Agrawala 2023) has been adapted for the video domain, facilitating the creation of structurally consistent frames in video generation (Khachatryan et al.
+2023) and translation tasks (Hu and Xu 2023; Zhang et al.
+2023b).
+Furthermore, GLIGEN’s (Li et al.
+2023) adaptation to the video domain by Ground-A-Video (Jeong and Ye 2023) demonstrates the use of both spatially-continuous depth map and spatially-discrete bounding box conditions, achieving multi-attribute editing of videos in a zero-shot manner.
+The advent of open-source text-to-video diffusion models (Wang et al.
+2023a; Sterling 2023; Chen et al.
+2023a; Wang et al.
+2023b) has spurred research into separating, altering, and combining the appearance and motion elements of videos (Zhao et al.
+2023b; Jeong, Park, and Ye 2023; Wei et al.
+2023; Bai et al.
+2024; Yang et al.
+2024; Yatim et al.
+2023; Ren et al.
+2024; Jeong et al.
+2024).
+MotionDirector (Zhao et al.
+2023b) and DreamVideo (Wei et al.
+2023) have each suggested approaches for dividing fine-tuning processes into distinct learning phases for subject appearance and temporal motion, utilizing efficient fine-tuning methods.
+On the other hand, VMC (Jeong, Park, and Ye 2023) focuses on distilling the motion within a video by calculating the residual vectors between consecutive frames.
 In their work, they fine-tune temporal attention layers within cascaded video diffusion models to synchronize the ground-truth motion vector with the denoised motion vector estimate, successfully generating videos that replicate the motion pattern of an input video within diverse visual scenarios.
-Similarly, [42] introduces a space-time feature loss that constructs self-similarity matrices based on the differences in attention features between frames.
+Similarly, (Yatim et al.
+2023) introduces a space-time feature loss that constructs self-similarity matrices based on the differences in attention features between frames.
 This approach aims to minimize the discrepancy in self-similarity between the input and output videos.
-0.B.2 Frequency-aware Visual Generation
+B.2 Frequency-aware Visual Generation
 Spectral analysis plays a pivotal role in the domain of visual understanding and generation, offering insights into the temporal-spatial structure of pixel-domain frames through frequency-domain signals.
-[23] introduced a hierarchical motion estimation algorithm employing complex discrete wavelet transforms, effectively utilizing phase differences among subbands to indicate local translations within video frames.
-[29] enhanced scalable video compression through motion-compensated wavelet transforms, integrating a continuous deformable mesh motion model to achieve superior compression efficiency and motion representation.
+(Magarey and Kingsbury 1998) introduced a hierarchical motion estimation algorithm employing complex discrete wavelet transforms, effectively utilizing phase differences among subbands to indicate local translations within video frames.
+(Secker and Taubman 2002) enhanced scalable video compression through motion-compensated wavelet transforms, integrating a continuous deformable mesh motion model to achieve superior compression efficiency and motion representation.
 Furthermore, these spectral insights have been instrumental in refining algorithms and deepening architectural understanding, particularly within the contexts of U-Net and autoencoder frameworks.
-[43] provided a groundbreaking reinterpretation of deep learning architectures for image reconstruction, establishing a connection between deep learning and classical signal processing theories, including wavelets and compressed sensing.
-[44] developed a wavelet-based correction method, WCT2, to augment photorealism in style transfer, leveraging whitening and coloring transforms to preserve structural integrity and statistics within the VGG feature space.
-In the realm of U-Net in diffusion models, [37] highlighted the rapid dominance of noise over high-frequency information in residual U-Net denoisers.
-Concurrently, [19] proposed the Hybrid Video Diffusion Model (HVDM), a novel architecture that captures spatiotemporal dependencies using a disentangled representation combining 2D projection and 3D convolutions with wavelet decomposition.
-While [19] aims to reconstruct input video with frequency matching loss, our approach does not aim to reconstruct input, focusing instead on learning motion dynamics through motion vectors, thereby distinguishing our method within the landscape of spectral analysis applications in motion estimation and transfer.
-Appendix 0.C Experimental details
+(Ye, Han, and Cha 2018) provided a groundbreaking reinterpretation of deep learning architectures for image reconstruction, establishing a connection between deep learning and classical signal processing theories, including wavelets and compressed sensing.
+(Yoo et al.
+2019) developed a wavelet-based correction method, WCT2, to augment photorealism in style transfer, leveraging whitening and coloring transforms to preserve structural integrity and statistics within the VGG feature space.
+In the realm of U-Net in diffusion models, (Williams et al.
+2024) highlighted the rapid dominance of noise over high-frequency information in residual U-Net denoisers.
+Concurrently, (Kim et al.
+2024) proposed the Hybrid Video Diffusion Model (HVDM), a novel architecture that captures spatiotemporal dependencies using a disentangled representation combining 2D projection and 3D convolutions with wavelet decomposition.
+While (Kim et al.
+2024) aims to reconstruct input video with frequency matching loss, our approach does not aim to reconstruct input, focusing instead on learning motion dynamics through motion vectors, thereby distinguishing our method within the landscape of spectral analysis applications in motion estimation and transfer.
+Appendix C Experimental details
 For spectral global motion alignment, we mainly use Haar wavelets with the number of levels $l=3$ for 8-frame videos and $l=4$ for 16-frame videos.
@@
 We follow other configurations, e.g.
-optimization algorithm, learning rate, training steps, etc., from the original motion transfer frameworks.
-Appendix 0.D Additional Results
+optimization algorithm, learning rate, training steps, etc., from the original motion transfer frameworks for a fair comparisons.
+Code is included as a supplementary material.
+Appendix D Additional analysis
+Ablation study.
+We first provide additional qualitative ablation results of global alignment in Fig.
+7 (baseline: VMC).
+Combined with quantiative ablation studies in the main paper, these demonstrates the effectiveness of both global and local motion alignment.
+Please refer to the Figure 6 in the main paper for more analysis on the local term.
+Training progress.
+To analyze the training progress further, we visualize the intermediate customization results with and without proposed spectral motion alignment.
+Notably, Fig.
+8a illustrates the gradual improvements of spatial layout attributed to the proposed global/local motion alignment.
+Specifically, our layout progressively aligns better as training proceeds.
+Both global and local terms contribute to this alignment as both match motion vectors $\delta\boldsymbol{v}_{0}$ encoding structural movement information.
+Alternative frequency-domain approaches.
+Our framework highlights two key insights: (a) susceptibility of 2D motion information in $\delta\boldsymbol{v}_{0}$ to spatial/motion artifacts, and (b) potential benefits of multi-scale decomposition for refining motion representation.
+In this context, both multi-scale Fourier and wavelet analyses offer theoretically viable options for spectral motion refinement.
+Demonstrating this, we use 2D Discrete Wavelet Transform (DWT) with Daubechies 3 wavelet, to obtain wavelet coefficients of ground truth and predicted motion vectors.
+Then each motion vector is reconstructed from these coefficients, excluding the finest high-frequency detail coefficients.
+Figure 8b illustrates that DWT-based refinement also significantly improves motion transfer compared to baseline (VMC), particularly by prioritizing low-frequency components of local motion encoding core motion information.
+While both methods show notable improvements, DFT outperforms DWT, with DWT requiring more fine-tuning of hyperparameters (e.g., the number of levels depending on spatial resolution, wavelet type, etc).
+Thus, we opt for DFT in refinement.
+Appendix E Additional Results
 This section provides additional qualitative comparisons of SMA across different baseline approaches.
-In Figure 8, we compare the efficacy of MotionDirector [49] in transferring motion, both with SMA integrated and without.
-Figures 9 and 10 delve into VMC’s [16] capabilities in customizing motion, showcasing outcomes with SMA and without its application.
-Additionally, Figure 11 contrasts the ability of Tune-A-Video and ControlVideo [38] to replicate the original motion, examining the impact of incorporating SMA.
-All our qualitative results can be viewed on our project page in video format.
-Figure 8: Additional comparison within MotionDirector framework.
+Figures 9 and 10, 11 delve into VMC’s (Jeong, Park, and Ye 2023) capabilities in customizing motion, showcasing outcomes with SMA and without its application.
+In Figure 12, we compare the efficacy of MotionDirector (Zhao et al.
+2023b) in transferring motion, both with SMA integrated and without.
+Additionally, Figure 13 contrasts the ability of ControlVideo (Wu et al.
+2023) to replicate the original motion, examining the impact of incorporating SMA.
+Figure 7: Ablation study on global alignment. Global motion alignment facilitates motion transfer.
+Figure 8: (a) Visualization of the training progress with and without local term (Global motion alignment leads to similar trends).
+(b) Comparisons of local motion refinement with 2D DWT and 2D DFT (originally proposed).
 Figure 9: Additional comparison within VMC framework, deployed on Show-1 Cascade.
 Figure 10: Additional comparison within VMC method, implemented on Zeroscope T2V.
-Figure 11: Additional comparison within Tune-A-Video and ControlVideo frameworks.
+Figure 11: Additional comparison within VMC method, implemented on Zeroscope T2V.
+Figure 12: Additional comparison within MotionDirector method.
+Figure 13: Additional comparison within ControlVideo frameworks.